### v1

# FrameFusion: Combining Similarity and Importance for Video Token Reduction on Large Visual Language Models

Abstract
[…]
1 Introduction
Large Vision-Language Models (LVLMs) have shown outstanding abilities across various video understanding scenarios, including temporal and spatial perception, recognition, and reasoning [34, 29, 14, 4]. Growing applications also demand LVLMs’ ability to process longer and more complex videos [21, 18, 24].
However, video understanding incorporates substantial computing overhead for LVLMs. Common LVLMs sample frames from the video, split each into image patches, and then sequentially embed them as vision tokens with a visual encoder. Though effective, it requires the processing of a huge amount of vision tokens for LVLMs. Take Google Gemini as an example, with the standard sampling rate of 1 frame per second (fps), it requires an astonishing amount of one million tokens to understand an hour’s video [21].
The importance-based token pruning has been extensively explored by previous works to solve the efficiency problem. They reduce vision tokens with various importance matrices, like accumulative attention score [23, 35, 7] and L2-norm of token feature [3]. However, there exist extensive important yet similar vision tokens in LVLMs due to frame resemblance and similar visual elements, while the effect of these tokens is less explored.
Figure 1: The central idea of FrameFusion. Compared with importance-based token pruning, FrameFusion additionally applies similarity-based token merging, keeping only important and unique vision tokens.
In this work, we recover similarity as a perpendicular property to importance for vision tokens reduction. The main idea is that although some frames are important for visual understanding, their high similarity leads to redundant information. These similar tokens, like those alike image patches from adjacent frames, hardly provide additional information, thus can be merged. To fully harness the potential of similarity-based merging, we study the token similarity distribution of LVLMs by answering three questions:
  1. 1.
    Where and when does similarity appear?
  2. 2.
    What is the value distribution of token similarity?
  3. 3.
    How consistent is the ranking of token similarity across layers?
Based on the observations, we propose FrameFusion to combine both similarity and importance for efficient LVLM token reduction. FrameFusion first evaluates the token similarity with minimum overhead and progressively merges similar tokens. After the similarity-based merging, FrameFusion then utilizes importance-based pruning to reduce tokens to the given computing budget.
Our contributions are three-folds:
  1. 1.
    We analyze the characteristics of token similarities across input positions and layers of LVLM.
  2. 2.
    We propose FrameFusion to hierarchically merge similar vision tokens and prune unimportant ones in a plug-and-play manner.
  3. 3.
    We conduct extensive experiments to show the effectiveness of FrameFusion on various video tasks.
Experiments show that FrameFusion pushes the Pareto front further for token compression. It successfully reduces the token budget and computing FLOPs to 30%, while maintaining comparable performance with the original dense model. The simple yet effective design helps FrameFusion to generalize across various model sizes, input lengths, and video tasks.
2 Related Work
2.1 Large Vision Language Model (LVLMs)
The LVLM architecture typically consists of a vision encoder and a Large Language Model (LLM) [34, 29, 14, 4, 18, 24, 13]. The vision encoder converts visual inputs into token sequences, which the LLM then processes alongside text sequences to generate responses. Specifically, for video input, frames are first sampled temporally and then specially divided into sequences of image patches before sending to the vision encoder [34, 14, 29], as shown in Figure 1. Due to the high temporal and spatial resolution demands of complex video understanding tasks, token lengths can reach up to one million for an hour-long video [21], imposing significant computational overhead on LVLMs.
2.2 Token Compression
Motivated by the heavy overhead of video processing, token compression becomes an essential method for LVLM efficiency. Existing methods compress tokens at three subsequent processes.
The first branch of work reduces the initial visual inputs before sending them to the vision encoder. They set rules to mix different temporal sampling frequencies [28, 34] and special resolutions [29, 5] when converting videos to input sequences, introducing trade-offs between visual detail and efficiency. Despite the simplicity, they inevitably incur direct detail losses and neglect the visual content guidance during compression.
Other works reduce tokens inside the vision encoder. They selectively retrieve [11] or condense [18] vision tokens in the vision encoder based on the guidance of text instructions. Yet, they require re-encoding all vision tokens if the text instruction changes, which incurs significant overheads for common multi-round conversation scenarios. Besides, an additional model fine-tuning is often needed to align the new vision encoding space.
Another branch of work focuses on token reduction in the subsequent LLM. For text-only tasks, previous works design static [8, 26, 10] or dynamic [12, 35, 9, 19] pruning pattern based on the importance of token (or KV-Cache). Emerging concurrent works highlight the specific token importance distribution for vision-language tasks [2, 16, 22, 33, 31], which further increase the sparsity of importance-based token pruning. However, as shown in Figure 4, token importance is inconsistent across different layers. It incurs prediction loss by pruning an unimportant token at shallow layers, which becomes important but inaccessible at deeper layers. FrameFusion falls in this category, exploring a more consistent and perpendicular token reduction method: similarity-based token merging.
3 Token Similarity Analysis
While token importance properties of LVLMs have been extensively studied [2, 16, 22, 33], a thorough and comparative analysis of token similarity remains lacking. Thus, we conduct oracle experiments to answer three key questions regarding the characteristics of token similarity, comparing them with token importance.
3.1 Experiment Setups and Notion Definitions
In this section, we present the oracle results with the Llava-Video-7B [34] model tested on 128 video samples (64 frames at 1 fps) from the comprehensive VideoMME [6] datasets. All metrics are averaged over all test samples unless otherwise noted. Results on different models are similar and provided in Appendix 10.2.
We define token importance and similarity using the notations $I^{(l)}\in\mathbb{R}^{N}$ and $S^{(l)}\in\mathbb{R}^{N}$, respectively, where $l$ indicate the layer index of the LLM. For clarity, we omit $l$ where it does not introduce ambiguity. We use the subscript $t$ to index the token along the input length dimension $N$. The token features at layer $l$ are represented as $\mathbf{X}^{l}\in\mathbb{R}^{N\times d}$.
Following previous works [2, 35], the token importance $I^{(l)}_{t}$ is quantified using the cumulative attention score, calculated by summing the post-softmax attention scores vertically across the $t$-th column and averaging this sum across all attention heads at layer $l$.
For token similarity $S^{(l)}_{t}$, we specifically define it as the cosine similarity between a vision token and its corresponding token from the preceding frame. The rationale for this definition is detailed in Section 3.2. Formally, if each frame is encoded into $P$ vision tokens, token similarity is calculated using the following equation:
$S_{t}=\frac{X_{t-P}^{T}X_{t}}{\lVert X_{t-P}\rVert_{2}\cdot\lVert X_{t}\rVert_{2}}$ (1)
3.2 Where and When Does Similarity Appear?
Figure 2: Token similarities among all input tokens at the first LVLM layer in Llava-Video-7B models. For visual clarity, the color bar displays only the top 90% of similarity values. Vision tokens begin at index 14, with 210 tokens per frame.
We first investigate which tokens are most likely to demonstrate high similarity, as these are expected to have a greater potential for merging. To ensure figure clarity within the limited resolution, we limit the number of frames so that each input sequentially consists of 14 system prompt tokens, 840 vision tokens (representing 4 frames at 210 tokens per frame), and 20 user instruction tokens. Figure 2 displays the $N\times N$ cosine similarity matrix for all input tokens at the first LLM layer. A prominent 210th sub-diagonal appears, reflecting high similarities between tokens $i$ and $i+P$, where $P=210$ corresponds to the number of vision tokens per frame. This pattern underscores our initial observation:
Observation 1. Spatially corresponding vision tokens between adjacent temporal frames exhibit higher cosine similarities than other token pairs.
Based on this observation, we focus on token similarities among these particular tokens and define token similarity as described in Equation 1. It is also worth noting that similar to token importance, the similarities between vision-language tokens are significantly lower than those between vision-vision or language-language tokens. This finding motivates a similar design choice to separate token compression for vision and language tokens, as proposed in prior work [2].
3.3 What Is the Distribution of Token Similarity?
Figure 3: Heatmap of token similarity across different model layers. Each cell represents a similarity range at a specific layer, with color intensity denoting distribution frequency. The line overlay shows the average token similarity at each layer.
To decide the appropriate way for token compression, we explore the distribution of token similarities across different layers. As illustrated in Figure 3, although the average token similarity across layers remains relatively stable, there is a significant shift in the distribution of these similarities:
Observation 2. High token similarities exhibit an obvious reduction in deeper model layers.
This trend is further confirmed by the variance in token similarity, which decreases from $6.0\times 10^{-2}$ at the first layer to $3.7\times 10^{-2}$ at the last layer. It illustrates a tightening of the distribution, with more numerical details in Appendix 10.1.
The causal attention mechanism in LLMs contributes to this shift, where later vision tokens can aggregate information from earlier tokens, but not vice versa. The attention range differences lead to a divergence in the similarity of tokens that were initially similar, such as corresponding tokens from adjacent frames. This divergence accumulates as layers deepen.
Based on this observation, FrameFusion prioritizes token merging at shallower layers rather than deeper layers to fully utilize the polarized high similarities.
3.4 How Consistent Is the Ranking of Token Similarity Across Layers?
The ranking consistency of token properties guides the compression strategy designs. For importance-based token pruning, the consistency across layers determines whether tokens deemed unimportant at shallow layers might become important at deeper layers. Early works [2, 23] adopt cascaded token pruning, where pruned tokens are no longer accessible in subsequent layers. These methods prioritize efficiency by eliminating token computations in both attention and Feed Forward Network (FFN) modules, and the importance is only computed once for pruned tokens. In contrast, more recent works [20, 17, 35] adopt non-cascaded pruning, which maintains the accessibility of all tokens across all layers by only pruning the KV-Cache at the attention layer. Yet, these methods do not reduce computation for the FFN and incur higher costs due to the need to compute importance for all tokens at every layer. Previous researches [8, 25] highlight the inconsistency of token importance across layers, which justifies re-evaluating pruning decisions at each layer. However, does a similar conclusion hold for token similarity?
Figure 4: Spearman Rank Correlation between adjacent layers for the Llava-Video-7B model.
To quantitatively assess the ranking consistency of token importance and token similarity, we calculate their respective Spearman Rank Correlations (SRC) [30] between adjacent layers. The SRC measures the degree to which the ranking of metrics in one layer correlates with the rankings in other layers. As shown in Figure 4, the SRCs for token similarity approach 1 across layers, indicating consistent ordering. In contrast, the SRCs for token importance are consistently lower, indicating weaker consistency across layers. We observe the following:
Observation 3. Token similarity demonstrates high rank consistency across different model layers.
Combining Observations 3 and 4, we conclude an interesting phenomenon: although the token similarity distribution condenses as layers deepen, tokens with the highest similarities remain the most similar. Motivated by this high consistency in similarity, FrameFusion employs similarity-based merging in a cascaded manner. Once highly-similar tokens are merged at shallow layers, they are not separated in subsequent computations.
Figure 5: The Top-30% retention rate across model layers using different retention metrics and starting layers.
We further validate this cascaded design choice by quantifying the top-k retention rate across model layers. For notion simplicity, we define token uniqueness as 1 minus similarity. To calculate the top-k retention rate, we identify the set of top-$k$ tokens under token uniqueness or importance metrics at each layer. The top-k retention rate for layer $l$, starting from layer $i$, is defined as the intersection rate between the top-k tokens from both layers $l$ and $i$.
To formalize the calculation of the top-k retention rate, we denote the set of indices of the top-$k$ retaining tokens based on the scoring function $f:\mathbb{R}^{N\times d}\rightarrow\mathbb{R}^{N}$ at layer $l$ as follows:
$T^{(l)}=\left\{i\;\big{|}\;i\in\operatorname{Top-K}(f(\mathbf{X}^{(l)}))\right\},$ (2)
where scoring function $f$ can be token importance or token similarity, $\operatorname{Top-K}:\mathbb{R}^{N}\rightarrow\mathbb{N}^{K}$ function returns the indices of the top-k values. The top-k retention rate $R^{(l)}_{i}$ of layer $l$ with the starting layer $i$ is calculated as:
$R^{(l)}_{i}=|T^{(l)}\cap T^{(i)}|/|T^{(i)}|.$ (3)
As shown in Figure 5, we quantify the top-30% retention rate with shallow starting layers 00 to $2$ under different metrics. The token similarity exhibits a much higher retention rate than token importance. This finding validates the superiority of using token similarity over importance for cascaded token pruning, achieving a better accuracy trade-off.
4 FrameFusion Design
Figure 6: FrameFusion first merges tokens with similarities above a specified threshold at shallow layers, then applies top-$k$ importance pruning to comply with the given computational constraints.
Building upon the observations from Section 3, we introduce FrameFusion, a novel token compression technique for video LVLMs, exploring a new perspective of token similarity. The design of FrameFusion is detailed in Section 4.1, followed by an explanation of the rationale behind each key design choice in Section 4.2.
4.1 Two-Stage Token Compression
The core concept of FrameFusion is illustrated in Figure 1: unlike traditional methods that primarily employ importance-based token pruning, FrameFusion also emphasizes similarity-based token merging. It merges similar tokens and prunes unimportant ones, retaining only those that are both important and unique. The two-stage token reduction process of FrameFusion is shown in Figure 6.
Merging stage. In the initial merging stage, FrameFusion utilizes token similarity to merge vision tokens. It computes the token similarity $S^{(l)}$ at the current layer using Equation 1, where only $N$ cosine similarities are computed between the corresponding vision tokens of adjacent frames. Tokens that exceed the similarity threshold $S_{\text{threshold}}$ are grouped with their analogous tokens from the previous frame. These merging groups are transitive, allowing concatenated groups to form a larger group containing more than two tokens as shown in Figure 6. Within each group, FrameFusion performs element-wise averaging of all tokens, assigning the result to the first token in the group. This forward merging strategy ensures that subsequent vision tokens can still aggregate information from all preceding tokens using causal attention.
The merging stage is applied across successive shallow LLM layers to progressively merge similar tokens, until the number of similar tokens falls below the threshold $N_{\text{threshold}}$ at a specific layer. Merging occurs right before the FFN module; however, to reduce the number of tokens for the attention module at the first LLM layer, an additional merging step is performed before it. After the merging stage, the remaining unique tokens advance to the pruning stage.
Pruning stage. After the merging stage, FrameFusion further prunes unimportant tokens. As discussed in Section 3.1, FrameFusion uses cumulative attention scores to represent token importance. Based on a user-defined computational cost budget, FrameFusion calculates the maximum number of remaining tokens $k$ that fits within the budget. It then applies top-$k$ importance pruning to retain only the important tokens from the remaining unique ones.
Through the merging and pruning stages, only the unique and important vision tokens are retained for subsequent processing, significantly boosting LVLM efficiency.
4.2 Design Choice Rationales
In this subsection, we explain the key design choices of FrameFusion with rationales from key observations in Section 3.
Unlike token importance, which reuses the existing $N\times N$ attention scores, token similarity introduces a new, orthogonal metric. To avoid additional $N\times N$ similarity computations for all token pairs, we leverage Observation 2 to compute only empirically similar token pairs with an $O(N)$ complexity:
Design Choice 1. FrameFusion computes token similarities only between corresponding vision tokens of adjacent frames.
Another key design choice is determining the layers at which to perform token reduction. For importance-based pruning, previous study [2] observes the reduction in vision token importance after the initial two layers for LVLMs. Thus, it advises applying importance-based pruning after the second layer. Other works also confirm the accuracy advantage of aggressive pruning only at deeper layers [8, 1]. In contrast, merging relies on high token similarities and therefore prefers shallow layers, according to Observation 3. Based on the contradictory preferences of similarity-based merging and importance-based pruning, FrameFusion is designed as follows:
Design Choice 2. FrameFusion applies token merging at the initial successive layers, then pruning at a later layer.
The final key design choice is whether merged tokens should remain reduced in subsequent layers, i.e., whether to use cascaded token merging. This choice depends on whether tokens merged in shallow layers are likely to be needed in deeper layers. Given the high rank consistency and retention rate of token similarities discussed in Observation 4, we adopt the following design to achieve a better efficiency trade-off:
Design Choice 3. FrameFusion merges tokens in a cascaded manner.
These rationales for the three key design choices in FrameFusion enhance the explainability of its better performance.
5 Experiment
5.1 Setups
Baselines. We compare FrameFusion with state-of-the-art token pruning baselines, StreamingLLM [26] and FastV [2]. Hyperparameters follow the official implementations and are detailed in Appendix7.
Token Budget. We term the token budget as the average number of tokens required for the KV-Cache at the start of the decoding stage. The relative token budget, denoted as $C$, is the token budget of compression methods divided by the original input length $N$. The relative token budget approximates the relative KV-Cache utilization for each compression method. We set $C=30\%$ for all token compression methods if not specified. Specifically, for the cascaded methods FrameFusion and FastV, the token budget equals the average token length per layer across both the prefill and decode stages. For StreamingLLM, the token budget equals the token length in KV-Cache at the decode stage, which also equals the attention span during the prefill stage.
Models. We evaluate our approach on widely used video LVLMs, including lmms-lab models: LLaVA-Video-{7B,72B}-Qwen2 (referred to as Llava-Video-{7B,72B}), LLaVA-NeXT-Video-32B-Qwen (referred to as Llava-Video-32B) [34], and the openbmb model MiniCPM-V-2_6 (referred to as MiniCPM-V-8B) [29].
Benchmarks. We use lmms-eval [32] as the primary evaluation framework to test diverse video benchmarks, including VideoMME [6] for video understanding, NExT-QA [27] for video question-answering, and Video Needle In A Haystack (VideoNIAH)[36] for visual content retrieval.
5.2 Accuracy-Computation Trade-off
Figure 7: The accuracy-computation trade-offs of various token compression methods, tested on Llava-Video-7B with VideoNIAH benchmark. Original* represents the original model with reduced frame rates.
As shown in Figure 7, we explore the accuracy-computation trade-off of different token compression methods. The relative FLOPs is normalized by the original dense model with the video sampling rate at 1 frame per second. For the original variant, the relative FLOPs is altered by directly multiplying the sampling rate with the relative token budget. FrameFusion advances the Pareto Front in video content retrieval accuracy and computing FLOPs across varied relative token budgets.
5.3 Performance
VideoMME $\uparrow$ NExt-QA $\uparrow$ VideoNIAH $\uparrow$
Model Method Short Medium Long Multi Choice Open End Edit Insert1 Insert2 Avg.
Llava-Video-7B Original 75.8 61.7 52.2 83.2 32.1 90.7 50.7 88.0 66.8
StreamingLLM 63.4 54.1 46.4 79.0 30.3 26.0 15.3 28.7 42.9
FastV 68.4 58.0 49.6 81.1 31.2 69.3 28.7 76.7 57.9
Ours 74.0 59.8 50.0 81.8 31.7 90.0 48.7 87.3 65.4
MiniCPM-V-8B Original 69.1 56.6 49.8 78.9 13.8 88.7 36.7 88.7 60.3
StreamingLLM 61.1 51.8 48.4 76.0 23.2 22.0 15.3 28.7 40.8
FastV 67.1 53.9 49.2 78.0 14.8 82.7 26.7 71.3 55.5
Ours 69.7 54.1 48.3 78.2 16.3 89.3 41.3 89.3 60.8
Table 1: The overall performance comparison across different models and methods on VideoMME, NextQA, and VideoNIAH benchmarks. All token compression methods employ 30% relative token budget.
FrameFusion outperforms state-of-the-art token compression methods across various models and benchmarks, achieving performance comparable to the dense model at a 30% token budget (Table 1). Additional results for additional budgets are in Appendix 8.1. Across VideoMME, NExT-QA, and VideoNIAH benchmarks, FrameFusion limits relative performance drops to under 3%. In contrast, StreamingLLM and FastV show drops of up to 13% and 7% for VideoMME and NExT-QA, and 70% and 24% for VideoNIAH. Further retrieval performance details are in Appendix 8.2.
5.4 Efficiency
Figure 8: Runtime and memory breakdown of Llava-Video-7B on a single A100-80GB GPU using FrameFusion. A relative token budget of 1.0 represents the original dense model. Numbers on bars show (a)LLM and end-to-end speedups and (b) LLM’s KV-Cache and total relative memory.
We evaluate efficiency on a single NVIDIA A100-80GB GPU, averaging results over 128 videos from the VideoMME benchmark. Figure 8 highlights how FrameFusion improves runtime and memory efficiency for the colored components.
Time. FrameFusion achieves 3.4-4.4$\times$ speedups with a 30% token budget across various input frames. Including other runtimes like ViT and CPU calls, FrameFusion maintains 1.6 – 1.9$\times$ end-to-end speedups. Speedup results for additional model sizes are provided in Appendix 8.3.
Memory. FrameFusion reduces LLM memory consumption for KV-Cache and corresponding activations (noted as “KV-Cache” in Figure 8) to 44% – 51% with a 30% token budget, enabling more frames to fit within GPU memory limits.
5.5 Scalability
Scaling Model Size.
Figure 9: The VideoMME performance for each category across Llava-Video-7B, 32B, and 72B for different methods. All scores are normalized by the original model.
We test the scalability of FrameFusion across diverse model parameter sizes with the Llava-Video model family. As shown in Figure 13, FrameFusion consistently outperforms FastV baseline across all model sizes and VideoMME categories, demonstrating comparative performance with the original model at a 30% relative token budget. The numeric results are shown in Appendix 8.5.
Figure 10: The VideoNIAH performance for the Llava-Video-7B across various numbers of input frames.
Scaling Input Length. Figure 10 shows that FrameFusion surpass other baselines to achieve high retrieval accuracies across diverse input number of frames, demonstrating comparable performance with the original dense model. Figure 8 shows increasing LLM speedups and memory reductions with more input frames.
5.6 Ablation Study
We validate the impact of each observation and design choice in FrameFusion, demonstrating their individual effectiveness in this section.
Figure 11: Top-k hit numbers for different similarity calculation strategies on the initial layers of the Llava-Video-7B model.
Design Choice 4.2. Instead of computing $N\times N$ token similarities with significant overhead, FrameFusion calculates $N$ token similarities only between corresponding vision tokens in adjacent frames. We compare our approach with two common alternatives: 1. The adjacent strategy, which computes $N$ similarities between adjacent image patches (i.e., adjacent vision tokens). 2. The random strategy, which calculates similarities for $N$ randomly selected token pairs. All three strategies are compared against the posterior optimal upper bound, which merges the most similar tokens using the full $N\times N$ similarity computation. As shown in Figure 11, we evaluate the top-K hit number for each strategy. This metric indicates how many of the top-k similar token pairs identified by a given strategy align with those selected by the posterior optimal approach. For visualization clarity, we limit the number of frames to three (420 non-first-frame vision tokens) from each video. Our method consistently outperforms the adjacent and random strategies across all model layers.
Merging Layer Budget Merging Rate Retrieval Acc.
Original 100% 0% 76.4%
0 50% 50.0% 76.2%
1 50% 52.0% 76.8%
2 50% 53.8% 76.4%
12 50% 87.5% 74.4%
13 50% 93.3% 64.2%
14 50% 100.0% 48.9%
Table 2: Performance on VideoNIAH of different merging layers with the same relative token budget of 50%.
Method VideoMME $\uparrow$ VideoNIAH $\uparrow$
Original 63.2 76.4
Prune then merge 59.9 73.1
Merge then prune 60.9 73.3
Table 3: Performance of the Llava-Video-7B model of different orders between merging and pruning.
Design Choice 4.2. FrameFusion first merges at the initial layers, then prunes at subsequent layers. We validate the influence of the merging layer position. As shown in Table 2, given the same token reduction budget, merging at shallow layers shows significantly better performance than merging at deeper layers. Additionally, we examine the effect of the order of merging and pruning. As shown in Table 3, with a fixed token budget of $C=0.3$ and the same number of tokens reduced in layers 1 and 2, merging before pruning achieves better performance compared to pruning before merging.
[…]
#Frame w. Cascaded Original Ours (Cascaded)
8 92.8% 20.0 20.7
16 90.0% 29.1 28.0
32 81.5% 43.8 42.9
64 72.2% 76.4 75.6
Table 4: Equivalent relative token budgets and VideoNIAH scores of cascaded merging across different numbers of frames, aligned to the minimum FLOPs achievable by non-cascaded merging.
Design Choice 4.2. FrameFusion adopts a cascaded merging strategy, where merged tokens remain reduced across layers to maximize efficiency. To evaluate the accuracy-efficiency trade-offs of cascaded merging, we use non-cascaded merging (KV-cache merging) as a baseline. This baseline reduces only the computation of attention matrices while leaving FFN computations unchanged. We calculate the minimum FLOPs achievable with non-cascaded merging by removing all KV-caches. As shown in Table 4, under the same FLOPs, FrameFusion requires token budgets of only 72% to 93% when using cascaded pruning—substantially more relaxed than the 30% budgets used in the main experiments. Across different numbers of frames, FrameFusion achieves performance comparable to the original model, while the non-cascaded model will fail due to the absence of KV-caches.
6 Conclusion
In this paper, we propose FrameFusion, a similarity-based token merging method for video LVLMs. By combining similarity-based merging with importance-based pruning, FrameFusion reduces redundant visual tokens while retaining critical information. This approach optimizes computational efficiency and memory usage, enabling accurate video understanding with significantly fewer tokens. Experiments across multiple benchmarks demonstrate a 70% reduction in vision tokens with minimal performance loss, achieving over 3.1$\times$ LLM speedups and 1.6$\times$ end-to-end speedups. FrameFusion providing new insights in token similarity for LVLMs, offering an efficient and scalable solution for real-world video language applications.
Supplementary Material
7 Detailed Experiment Setup
For the baselines StreamingLLM [26] and FastV [2], we follow the official implementations and set the attention sink size of StreamingLLM to 8 and $K$ in FastV to 2.
For FrameFusion, the merging ratios across layers are controlled by two hyperparameters: $S_{\text{threshold}}$ and $N_{\text{threshold}}$, as discussed in Section 4.1.
$S_{\text{threshold}}$ defines the minimum cosine similarity required for two tokens to be considered similar and merged. Since similarity distributions vary across models, we set $S_{\text{threshold}}$ to match the median of similarity at the first model layer under typical input cases, such as 128 samples from the VideoMME dataset. For the Llava-Video series, we set $S_{\text{threshold}}=0.6$; while for MiniCPM-V, we set $S_{\text{threshold}}=0.7$.
$N_{\text{threshold}}$ determines the transition from merging to pruning. If the number of similar tokens (tokens with cosine similarity above $S_{\text{threshold}}$) falls below $N_{\text{threshold}}$, the model switches to pruning. We set $N_{\text{threshold}}=0.1$ to avoid extensive similarity computations across the entire model.
To ensure the merging process does not excessively reduce the token count below the predefined token budget $C$, we precompute the maximum number of token pairs ($N_{\text{max}}$) that can be merged per layer. If the actual number of pairs exceeds $N_{\text{max}}$, only the top $N_{\text{max}}$ pairs with the highest cosine similarity are merged. Any remaining merging or pruning steps are skipped, and the model proceeds with a standard forward pass.
8 Additional Experiment Results
8.1 Performance Across Token Budgets
VideoMME NExt-QA-MC NExt-QA-OE
Model Method Budget Score $\uparrow$ Drop $\downarrow$ Score $\uparrow$ Drop $\downarrow$ Score $\uparrow$ Drop $\downarrow$ Max. Drop $\downarrow$
Llava-Video-7B Original 1.0 63.2 - 83.2 - 32.1 - -
Ours 0.3 61.3 3.0% 81.8 1.7% 31.7 1.2% 3.0%
0.5 62.6 0.9% 82.7 0.6% 32.1 0.0% 0.9%
0.7 63.0 0.3% 82.8 0.5% 32.1 0.0% 0.5%
MiniCPM-V-8B Original 1.0 58.5 - 78.9 - 13.8 - -
Ours 0.3 57.4 1.9% 78.2 0.9% 16.3 -18.1% 1.9%
0.5 58.5 0.0% 78.6 0.4% 17.4 -26.1% 0.4%
0.7 57.8 1.2% 78.6 0.4% 16.1 -16.7% 1.2%
Table 5: Performance comparison between the original and proposed methods on VideoMME, NExt-QA-MC, and NExt-QA-OE benchmarks with different relative token budgets on Llava-Video-7B model. Drop indicates the relative performance decrease compared to the original method.
Table 5 presents the benchmark performance of the Llava-Video-7B model at token budgets ranging from 0.3 to 0.7. At a 30% token budget, FrameFusion achieves strong performance, with a maximum relative drop of less than 3.0% compared to the dense model. As the budget increases to 0.5 and 0.7, the maximum drops further decrease to $\leq$1.2%.
8.2 Retrieval Analysis
Figure 12: VideoNIAH retrieval accuracy of the Llava-Video-7B and MiniCPM-V-8B models using different token compression methods across varying video lengths and retrieval positions. All token compression methods employ 30% relative token budget.
We further investigate the retrieval accuracy details with the VideoNIAH benchmark, as shown in Figure 12. FrameFusion demonstrates similar retrieval performance as the original dense model, with consistent performance across lengths and positions. In contrast, StreamingLLM hardly retrieves the initial frames of the video. FastV does not show particular failure patterns but undergoes uniform performance degradation across grids.
8.3 Speedup Across Model Sizes
We evaluate the scalability of FrameFusion ’s speedup across different model sizes, as shown in Figure 13. To accommodate the increased KV-Cache and memory overhead, we distribute the models across multiple GPUs. With larger models, FrameFusion achieves greater end-to-end speedups, delivering 2.8×2.8$\times$2.8 × for Llava-Video-32B on two GPUs and 3.2×3.2$\times$3.2 × for Llava-Video-72B on four GPUs at a 30% token budget.
Figure 13: Runtime of (a) Llava-Video-32B on two A100-80GB GPUs and (b) Llava-Video-72B on four A100-80GB GPUs using FrameFusion. Numbers on bars indicate LLM and end-to-end speedups.
8.4 Token Reduction Details
Figure 14: Average number of tokens per layer in the Llava-Video-7B model with FrameFusion at different relative token budgets. Error bars represent variance across data items.
FrameFusion reduces computational cost through both token merging and pruning. Using 128 samples from the VideoMME dataset with the Llava-Video-7B model, we calculate the token count per layer. As shown in Figure 14, FrameFusion progressively reduces tokens per layer, achieving the desired relative token budget (represented by the area under the line).
8.5 Scalability Numeric Results
We present the detailed numeric results of the scalability experiments in Section 5.5.
Model Method Knowledge Film & Television Sports Competition Artistic Performance Life Record Multilingual
Llava-Video-7B Original 63.1 67.2 61.8 61.7 63.7 58.9
StreamingLLM 55.1 57.2 56.0 54.2 52.9 48.9
FastV 59.1 60.0 58.9 57.8 58.1 55.6
Ours 62.7 63.6 58.0 61.7 60.8 56.7
Llava-Video-32B Original 60.9 68.9 57.3 59.2 58.1 58.9
StreamingLLM 53.1 55.6 50.2 53.3 49.5 53.3
FastV 56.1 62.2 55.6 56.7 54.3 50.0
Ours 58.9 65.6 55.6 59.2 54.3 54.4
Llava-Video-72B Original 73.2 74.4 68.0 71.4 68.9 62.2
StreamingLLM 65.7 66.7 59.3 65.6 58.7 58.9
FastV 66.8 72.8 61.1 69.2 63.2 61.1
Ours 72.2 73.1 65.6 69.7 65.9 61.1
Table 6: Numeric VideoMME scores of different methods and model sizes across various video categories.
Table 6 provides the VideoMME scores for various model sizes across different video categories, offering a numerical breakdown of Figure 13. FrameFusion consistently outperforms the baseline methods across diverse model sizes and video categories.
Number of frames Max.
Method 64 85 107 128 Relative Drop
Original 76.4 78.4 80.7 82.9 -
StreamingLLM 23.3 25.8 27.6 27.6 70%
FastV 58.2 63.6 65.8 69.3 24%
Ours 75.3 78.2 80.0 83.6 1%
Table 7: Numeric VideoNIAH retrieval accuracy of different methods across various frame counts.
Table 7 illustrates how retrieval accuracy scales with the number of input frames, complementing the insights from Figure 10. As shown, FrameFusion maintains consistent accuracy improvements across increasing frame numbers, matching the performance of the original model. In contrast, both StreamingLLM and FastV exhibit noticeable drops in accuracy.
9 Asymptotic Complexity Analysis
We estimate the computing cost of FrameFusion following the approach of FastV [2]. Given a model with $L$ layers and a specified relative token budget $C$, FrameFusion operates in the merging stage from layer 0 to layer $K-1$, then transitions to the pruning stage at layer $K$. Let $N_{l}$ denote the number of tokens layer $l$ before token reduction at this layer. Note that $N_{l+1}$ represents the number of tokens of layer $l$ after token reduction, and we let $N_{-1}$ equal the original input token length $N$. FrameFusion reduces $N_{l}$ with merging and pruning at the initial $K+1$ layers. After the token reduction, the remaining tokens for the successive layers are calculated as follows:
$N_{l}=\frac{L\times C\times N-(N_{0}+\ldots+N_{K})}{L-K-1},l\in[K+1,L)$ (4)
The model inference computation FLOPs $F(N_{l},N_{l+1})$ of layer $l$ is calculated as follows:
$F(N_{l},N_{l+1})=4N_{l}D^{2}+2N_{l}^{2}D+3N_{l+1}DM$ (5)
where $D$ denotes the hidden state size, and $M$ denotes the intermediate FFN size. The additional computation $F^{\prime}(N_{l})$ introduced by FrameFusion during similarity computation is:
$F^{\prime}(N_{l})=3N_{l}D$ (6)
Note that the additional computation $F^{\prime}$ introduced by FrameFusion shows neglectable asymptotic complexity with respect to input length and model size, comparing with the $O(N^{2}D)$ and $O(ND^{2})$ complexities of the original model.
10 Additional Observation Details
10.1 Similarity Distribution Details
Figure 15: Average token similarity variance per LLM layer in the Llava-Video-7B model, tested on 128 samples from the VideoMME dataset. Shading represents the variance across data items.
We take 128 videos from the VideoMME dataset and calculate the variance in token similarity across different layers. As shown in Figure 15, the similarity variance decreases in the deeper layers of the model, validating Observation 3.
10.2 Observations on Additional Models
In addition to the analysis of the Llava-Video model in Section 3, we conduct a similar study on the MiniCPM architecture. Results are presented in Figures 16, 17, 18, and 19.
Overall, the conclusions align with those of the Llava-Video model, with a few notable differences: Firstly, as shown in Figure 16, MiniCPM, which incorporates Q-Former [29, 15], exhibits additional high similarity among vision tokens within the same frame. However, the prominent 210th sub-diagonal persists, supporting our token similarity calculation strategy. Secondly, as shown in Figure 17, high similarity decreases less steeply in deeper layers for MiniCPM compared to Llava-Video. Despite this, the superior efficiency of cascaded merging at shallower layers ensures that Design Choice 4.2 remains valid.
Figure 16: Token similarities between all input tokens at the first LVLM layer in MiniCPM-V-8B.
Figure 17: Heatmap of token similarity across different model layers for the MiniCPM-V-8B model. Each cell represents the similarity at a specific layer, with color intensity denoting distribution frequency. The line overlay shows the average token similarity across layers.
Figure 18: Spearman Rank Correlation between adjacent layers for the MiniCPM-V-8B model.
Figure 19: The Top-30% retention rate across model layers for the MiniCPM-V-8B model, using different retention metrics and reference layers.
10.3 Video Pruning Visualization
Frame 0
Frame 1
Frame 2
Frame 3
Frame 4
Frame 5
Frame 6
Frame 7
Frame 8
Frame 9
Frame 10
Frame 11
Frame 12
Frame 13
Frame 14
Frame 15
Frame 16
Frame 17
Frame 18
Frame 19
Frame 20
Frame 21
Frame 22
Frame 23
Frame 24
Frame 25
Frame 26
Frame 27
Frame 28
Frame 29
Frame 30
Frame 31
Frame 32
Frame 33
Frame 34
Frame 35
Frame 36
Frame 37
Frame 38
Frame 39
Frame 40
Frame 41
Frame 42
Frame 43
Frame 44
Frame 45
Frame 46
Frame 47
Frame 48
Frame 49
Frame 50
Frame 51
Frame 52
Frame 53
Frame 54
Frame 55
Frame 56
Frame 57
Frame 58
Frame 59
Frame 60
Frame 61
Frame 62
Frame 63
Figure 20: An example input video with 1 fps frame rate.
Frame 0
Frame 1
Frame 2
Frame 3
Frame 4
Frame 5
Frame 6
Frame 7
Frame 8
Frame 9
Frame 10
Frame 11
Frame 12
Frame 13
Frame 14
Frame 15
Frame 16
Frame 17
Frame 18
Frame 19
Frame 20
Frame 21
Frame 22
Frame 23
Frame 24
Frame 25
Frame 26
Frame 27
Frame 28
Frame 29
Frame 30
Frame 31
Frame 32
Frame 33
Frame 34
Frame 35
Frame 36
Frame 37
Frame 38
Frame 39
Frame 40
Frame 41
Frame 42
Frame 43
Frame 44
Frame 45
Frame 46
Frame 47
Frame 48
Frame 49
Frame 50
Frame 51
Frame 52
Frame 53
Frame 54
Frame 55
Frame 56
Frame 57
Frame 58
Frame 59
Frame 60
Frame 61
Frame 62
Frame 63
Figure 21: The example of the video after token merging. Merged tokens are visualized with the blank blocks.
Frame 0
Frame 1
Frame 2
Frame 3
Frame 4
Frame 5
Frame 6
Frame 7
Frame 8
Frame 9
Frame 10
Frame 11
Frame 12
Frame 13
Frame 14
Frame 15
Frame 16
Frame 17
Frame 18
Frame 19
Frame 20
Frame 21
Frame 22
Frame 23
Frame 24
Frame 25
Frame 26
Frame 27
Frame 28
Frame 29
Frame 30
Frame 31
Frame 32
Frame 33
Frame 34
Frame 35
Frame 36
Frame 37
Frame 38
Frame 39
Frame 40
Frame 41
Frame 42
Frame 43
Frame 44
Frame 45
Frame 46
Frame 47
Frame 48
Frame 49
Frame 50
Frame 51
Frame 52
Frame 53
Frame 54
Frame 55
Frame 56
Frame 57
Frame 58
Frame 59
Frame 60
Frame 61
Frame 62
Frame 63
Figure 22: The example of the video after token merging. Merged tokens are visualized with the average image patches.
We select a video example to visualize the effect of our token merging strategy. Figure 20 shows the frames of the original video sampled at a frame rate of 1 fps. In Figure 21, we present the video input to the model after token merging in Layer 0, where blank patches indicate tokens that have been merged. Furthermore, we replace the blank regions with the average of the merged patches, and the resulting visualization is shown in Figure 22. As shown in the examples, FrameFusion token merging strategy successfully merges similar vision tokens, reducing the computational costs, while maintaining high vadility of the video.
10.4 Importance-Similarity Joint-Distribution
We visualize the joint distribution of token importance and similarity across different layers of Llava-Video-7B. As shown in Figure 23, it can be observed that in the shallow layers of the model, a significant number of tokens exhibit both high similarity and high importance values. FrameFusion can effectively compress these tokens. This phenomenon becomes less apparent in the deeper layers of the model, supporting our design choice of performing token merging in the shallow layers of the model.
(a) Layer 0
(b) Layer 1
(c) Layer 2
(d) Layer 14
(e) Layer 15
(f) Layer 16
(g) Layer 25
(h) Layer 26
(i) Layer 27
Figure 23: Importance-similarity joint-distribution of different layers, with color intensity denoting distribution frequency.
11 Limitation and Future Works
While FrameFusion demonstrates significant improvements in token reduction and efficiency for video LVLMs, certain challenges remain for future works. First, the similarity-based merging process can be further refined to better handle highly diverse or complex video content, minimizing potential information loss. Second, the reliance on pre-defined similarity and importance metrics calls for the development of adaptive and task-specific strategies to improve generalization across diverse scenarios. Future work will focus on designing more robust similarity measures and integrating FrameFusion with advanced token-efficient architectures.

### v2

@@
-FrameFusion: Combining Similarity and Importance for Video Token Reduction on Large Visual Language Models
+FrameFusion: Combining Similarity and Importance for Video Token Reduction on Large Vision Language Models
 Abstract
-[…]
+The increasing demand to process long and high-resolution videos significantly burdens Large Vision-Language Models (LVLMs) due to the enormous number of visual tokens.
+Existing token reduction methods primarily prune tokens based on importance metrics, such as cumulative attention scores.
+However, even important tokens may exhibit high redundancy caused by similarity among adjacent video frames and repetitive visual elements.
+To address this limitation, we propose FrameFusion, a novel token reduction approach integrating similarity-based merging with importance-based pruning.
+We conduct a thorough study on token similarity characteristics, revealing three key insights: (1) spatially corresponding visual tokens between adjacent frames have higher cosine similarities compared to other token pairs; (2) high token similarities prominently decrease in deeper model layers; and (3) token similarity rankings are highly consistent across different layers.
+Guided by these observations, FrameFusion computes token similarities exclusively between corresponding visual tokens from adjacent frames, applies token merging at initial successive layers followed by pruning in deeper layers, and adopts a cascaded merging strategy to further enhance efficiency.
+We evaluate FrameFusion comprehensively across six diverse LVLMs, ranging from 2B to 72B parameters, using five video benchmarks encompassing video retrieval, question-answering, and spatial-temporal understanding tasks.
+Experiments show that FrameFusion reduces visual tokens by 70%, achieving 1.6 – 3.6$\times$ end-to-end speedups, with an average performance impact of less than 3%.
+Our code is available at https://github.com/thu-nics/FrameFusion.
 1 Introduction
-Large Vision-Language Models (LVLMs) have shown outstanding abilities across various video understanding scenarios, including temporal and spatial perception, recognition, and reasoning [34, 29, 14, 4].
-Growing applications also demand LVLMs’ ability to process longer and more complex videos [21, 18, 24].
-However, video understanding incorporates substantial computing overhead for LVLMs.
-Common LVLMs sample frames from the video, split each into image patches, and then sequentially embed them as vision tokens with a visual encoder.
-Though effective, it requires the processing of a huge amount of vision tokens for LVLMs.
-Take Google Gemini as an example, with the standard sampling rate of 1 frame per second (fps), it requires an astonishing amount of one million tokens to understand an hour’s video [21].
-The importance-based token pruning has been extensively explored by previous works to solve the efficiency problem.
-They reduce vision tokens with various importance matrices, like accumulative attention score [23, 35, 7] and L2-norm of token feature [3].
-However, there exist extensive important yet similar vision tokens in LVLMs due to frame resemblance and similar visual elements, while the effect of these tokens is less explored.
+Large Vision-Language Models (LVLMs) have demonstrated remarkable capabilities across various video understanding tasks, including temporal and spatial perception, recognition, and reasoning [42, 34, 16, 5].
+Increasingly demanding applications require LVLMs to process longer and more complex videos [27, 2, 21, 30].
+However, handling extensive video data incurs substantial computational overhead.
+Typically, LVLMs sample frames from videos, divide each frame into image patches, and embed these sequentially as visual tokens through a visual encoder.
+While effective, this process generates an enormous number of tokens.
+For instance, Google’s Gemini, with a standard sampling rate of 1 frame per second (fps), needs to process approximately one million tokens to analyze an hour-long video [27].
 Figure 1: The central idea of FrameFusion.
-Compared with importance-based token pruning, FrameFusion additionally applies similarity-based token merging, keeping only important and unique vision tokens.
-In this work, we recover similarity as a perpendicular property to importance for vision tokens reduction.
-The main idea is that although some frames are important for visual understanding, their high similarity leads to redundant information.
-These similar tokens, like those alike image patches from adjacent frames, hardly provide additional information, thus can be merged.
-To fully harness the potential of similarity-based merging, we study the token similarity distribution of LVLMs by answering three questions:
+Compared with importance-based token pruning, FrameFusion additionally applies similarity-based token merging, keeping only important and unique visual tokens.
+Previous works primarily employ importance-based token pruning methods to mitigate efficiency demands.
+These approaches reduce visual tokens based on metrics such as cumulative attention scores [29, 43, 8] or normalized token feature [4].
+Nevertheless, among the top 10% of tokens ranked by cumulative attention scores, 55% of them exhibit high redundancy, with a cosine similarity above 0.9 in the first layer of Llava-Video-7B.
+In this work, we revisit token similarity as a perpendicular factor to token importance for reducing visual tokens.
+We argue that even important tokens can introduce redundancy due to visual similarities, particularly among adjacent frames.
+By merging these highly similar tokens, redundancy can be significantly reduced without compromising essential visual information.
+To effectively utilize token similarity, we first systematically investigate its characteristics within LVLMs.
+We find that (1) token similarity predominantly occurs between spatially corresponding tokens from adjacent frames, (2) similarities exhibit high values particularly at shallow layers, and (3) the token similarity rankings are highly consistent across layers.
+Based on these insight, we propose FrameFusion, a plug-and-play approach that integrates similarity-based merging with importance-based pruning.
+FrameFusion efficiently computes token similarities, progressively merges similar tokens at shallow layers, and subsequently applies importance-based pruning to adhere to computational constraints.
+Our contributions are summarized as follows:
   1. 1.
-    Where and when does similarity appear?
+    We systematically analyze token similarity characteristics across input positions and layers in LVLMs.
   2. 2.
-    What is the value distribution of token similarity?
+    We propose FrameFusion, a novel, post-training method that integrates similarity-based token merging with importance-based pruning for video token reduction.
   3. 3.
-    How consistent is the ranking of token similarity across layers?
-Based on the observations, we propose FrameFusion to combine both similarity and importance for efficient LVLM token reduction.
-FrameFusion first evaluates the token similarity with minimum overhead and progressively merges similar tokens.
-After the similarity-based merging, FrameFusion then utilizes importance-based pruning to reduce tokens to the given computing budget.
-Our contributions are three-folds:
-  1. 1.
-    We analyze the characteristics of token similarities across input positions and layers of LVLM.
-  2. 2.
-    We propose FrameFusion to hierarchically merge similar vision tokens and prune unimportant ones in a plug-and-play manner.
-  3. 3.
-    We conduct extensive experiments to show the effectiveness of FrameFusion on various video tasks.
-Experiments show that FrameFusion pushes the Pareto front further for token compression.
-It successfully reduces the token budget and computing FLOPs to 30%, while maintaining comparable performance with the original dense model.
-The simple yet effective design helps FrameFusion to generalize across various model sizes, input lengths, and video tasks.
+    We validate the effectiveness of FrameFusion through extensive experiments across diverse LVLMs, model sizes, input lengths, and video benchmarks.
+Experiments confirm that FrameFusion effectively advances the Pareto front in token compression.
+It reduces visual tokens to 30%, achieving 1.6–3.6$\times$ end-to-end speedup while maintaining less than a 3% performance drop over dense models.
+Its simple and effective design ensures broad applicability across various tasks and scenarios.
 2 Related Work
 2.1 Large Vision Language Model (LVLMs)
-The LVLM architecture typically consists of a vision encoder and a Large Language Model (LLM) [34, 29, 14, 4, 18, 24, 13].
-The vision encoder converts visual inputs into token sequences, which the LLM then processes alongside text sequences to generate responses.
-Specifically, for video input, frames are first sampled temporally and then specially divided into sequences of image patches before sending to the vision encoder [34, 14, 29], as shown in Figure 1.
+The LVLM architecture typically consists of a visual encoder and a Large Language Model (LLM) [42, 34, 16, 5, 21, 30, 15].
+The visual encoder converts visual inputs into token sequences, which the LLM then processes alongside text sequences to generate responses.
+Specifically, for video input, frames are first sampled temporally and then specially divided into sequences of image patches before sending to the visual encoder [42, 16, 34], as shown in Figure 1.
 Due to the high temporal and spatial resolution demands of complex video understanding tasks, token lengths can reach up to one million for an hour-long video [27], imposing significant computational overhead on LVLMs.
@@
 Motivated by the heavy overhead of video processing, token compression becomes an essential method for LVLM efficiency. Existing methods compress tokens at three subsequent processes.
-The first branch of work reduces the initial visual inputs before sending them to the vision encoder.
+The first branch of work reduces the initial input before sending them to visual encoder.
 They set rules to mix different temporal sampling frequencies [33, 42] and special resolutions [34, 6] when converting videos to input sequences, introducing trade-offs between visual detail and efficiency.
-Despite the simplicity, they inevitably incur direct detail losses and neglect the visual content guidance during compression.
-Other works reduce tokens inside the vision encoder.
-They selectively retrieve [11] or condense [18] vision tokens in the vision encoder based on the guidance of text instructions.
-Yet, they require re-encoding all vision tokens if the text instruction changes, which incurs significant overheads for common multi-round conversation scenarios.
-Besides, an additional model fine-tuning is often needed to align the new vision encoding space.
+Despite simplicity, they inevitably incur direct detail losses and neglect the content guidance for compression.
+Other works reduce tokens inside the visual encoder.
+They selectively retrieve [12] or condense [21] visual tokens in the visual encoder.
+Yet, they require re-encoding all visual tokens if the text instruction changes, which incurs significant overheads for common multi-round conversation scenarios.
+Besides, an additional model fine-tuning is often needed to align the new visual encoding space [13].
 Another branch of work focuses on token reduction in the subsequent LLM.
@@
 FrameFusion falls in this category, exploring a more consistent and perpendicular token reduction method: similarity-based token merging.
+More related works are discussed in Appendix 11.
 3 Token Similarity Analysis
-While token importance properties of LVLMs have been extensively studied [2, 16, 22, 33], a thorough and comparative analysis of token similarity remains lacking.
-Thus, we conduct oracle experiments to answer three key questions regarding the characteristics of token similarity, comparing them with token importance.
-3.1 Experiment Setups and Notion Definitions
-In this section, we present the oracle results with the Llava-Video-7B [34] model tested on 128 video samples (64 frames at 1 fps) from the comprehensive VideoMME [6] datasets.
-All metrics are averaged over all test samples unless otherwise noted.
-Results on different models are similar and provided in Appendix 10.2.
-We define token importance and similarity using the notations $I^{(l)}\in\mathbb{R}^{N}$ and $S^{(l)}\in\mathbb{R}^{N}$, respectively, where $l$ indicate the layer index of the LLM.
-For clarity, we omit $l$ where it does not introduce ambiguity.
-We use the subscript $t$ to index the token along the input length dimension $N$.
-The token features at layer $l$ are represented as $\mathbf{X}^{l}\in\mathbb{R}^{N\times d}$.
-Following previous works [2, 35], the token importance $I^{(l)}_{t}$ is quantified using the cumulative attention score, calculated by summing the post-softmax attention scores vertically across the $t$-th column and averaging this sum across all attention heads at layer $l$.
-For token similarity $S^{(l)}_{t}$, we specifically define it as the cosine similarity between a vision token and its corresponding token from the preceding frame.
-The rationale for this definition is detailed in Section 3.2.
-Formally, if each frame is encoded into $P$ vision tokens, token similarity is calculated using the following equation:
-$S_{t}=\frac{X_{t-P}^{T}X_{t}}{\lVert X_{t-P}\rVert_{2}\cdot\lVert X_{t}\rVert_{2}}$ (1)
-3.2 Where and When Does Similarity Appear?
+While token importance in LVLMs has been extensively explored [3, 18, 28, 41, 39], the characteristics of token similarity remain under-investigated.
+To bridge this gap, we conduct comprehensive oracle experiments analyzing token similarity and contrasting it with token importance.
+3.1 Experimental Setup and Definitions
+In this section, we present oracle experiment results on the Llava-Video-7B [42] model using the first 128 video samples from the VideoMME dataset [7], each comprising 64 frames sampled at 1 fps.
+Metrics reported are averaged over all samples unless otherwise noted.
+Similar results on other models are included in Appendix 10.
+We define token importance and token similarity as $I^{(l)}\in\mathbb{R}^{N}$ and $S^{(l)}\in\mathbb{R}^{N}$, respectively, with $l$ indicating the LLM layer index and $N$ denoting the number of input tokens.
+We use subscript $t$ to index the token along the input length dimension $N$.
+For simplicity, we omit the layer index when contextually clear.
+Input hidden features for layer $l$ are denoted as $\mathbf{X}^{(l)}\in\mathbb{R}^{N\times d}$.
+Following previous works [3, 43], token importance $I^{(l)}_{t}$ is computed using the cumulative attention score, calculated by summing the post-softmax attention scores vertically across the $t$-th column and averaging this sum across all attention heads at layer $l$.
+For token similarity $S_{t}^{(l)}$, we define it specifically between each visual token and its spatially corresponding token from the preceding frame, based on our empirical observation detailed in Section 3.2.
+Formally, token similarity is computed with cosine similarity:
+$S_{t}=\frac{X_{t-P}^{T}X_{t}}{\lVert X_{t-P}\rVert_{2}\cdot\lVert X_{t}\rVert_{2}},$ (1)
+where $P$ represents the number of visual tokens per frame, and $T$ represents the matrix transpose.
+3.2 Where Does High Similarity Occur?
 Figure 2: Token similarities among all input tokens at the first LVLM layer in Llava-Video-7B models.
 For visual clarity, the color bar displays only the top 90% of similarity values.
-Vision tokens begin at index 14, with 210 tokens per frame.
-We first investigate which tokens are most likely to demonstrate high similarity, as these are expected to have a greater potential for merging.
-To ensure figure clarity within the limited resolution, we limit the number of frames so that each input sequentially consists of 14 system prompt tokens, 840 vision tokens (representing 4 frames at 210 tokens per frame), and 20 user instruction tokens.
-Figure 2 displays the $N\times N$ cosine similarity matrix for all input tokens at the first LLM layer.
-A prominent 210th sub-diagonal appears, reflecting high similarities between tokens $i$ and $i+P$, where $P=210$ corresponds to the number of vision tokens per frame.
-This pattern underscores our initial observation:
-Observation 1. Spatially corresponding vision tokens between adjacent temporal frames exhibit higher cosine similarities than other token pairs.
-Based on this observation, we focus on token similarities among these particular tokens and define token similarity as described in Equation 1.
-It is also worth noting that similar to token importance, the similarities between vision-language tokens are significantly lower than those between vision-vision or language-language tokens.
-This finding motivates a similar design choice to separate token compression for vision and language tokens, as proposed in prior work [2].
-3.3 What Is the Distribution of Token Similarity?
-Figure 3: Heatmap of token similarity across different model layers.
+Visual tokens begin at index 14, with 210 tokens per frame.
+We first analyze which tokens typically exhibit high similarity, as these tokens have greater potential for merging.
+To maintain visual clarity, we limit the video frames to 4 (resulting in 840 visual tokens at $P=210$ tokens per frame), with 14 system prompt tokens and 20 user instruction tokens before and after the visual tokens.
+Figure 2 illustrates the $N\times N$ cosine similarity matrix at the first LVLM layer.
+A distinct 210th sub-diagonal emerges, highlighting significantly higher similarities between tokens at positions $i$ and $i+P$.
+Statistically, the average cosine similarity of 0.62 between these tokens far exceeds the average similarity (0.28) observed elsewhere.
+Hence, we conclude:
+Observation 1. Spatially corresponding visual tokens from adjacent frames exhibit higher cosine similarity compared to other token pairs.
+Based on this observation, we focus on similarities among these particular tokens and define token similarity as described in Equation 1.
+Additionally, sub-diagonals at multiples of 210 tokens also show relatively high similarities.
+This indicates that visual redundancy extends across multiple consecutive frames, which can also be identified by sequentially examining adjacent frames.
+3.3 What Is the Token Similarity Distribution Across Layers?
+Figure 3: Heatmap of token similarity across model layers.
 Each cell represents a similarity range at a specific layer, with color intensity denoting distribution frequency.
-The line overlay shows the average token similarity at each layer.
-To decide the appropriate way for token compression, we explore the distribution of token similarities across different layers.
-As illustrated in Figure 3, although the average token similarity across layers remains relatively stable, there is a significant shift in the distribution of these similarities:
-Observation 2. High token similarities exhibit an obvious reduction in deeper model layers.
-This trend is further confirmed by the variance in token similarity, which decreases from $6.0\times 10^{-2}$ at the first layer to $3.7\times 10^{-2}$ at the last layer.
-It illustrates a tightening of the distribution, with more numerical details in Appendix 10.1.
-The causal attention mechanism in LLMs contributes to this shift, where later vision tokens can aggregate information from earlier tokens, but not vice versa.
-The attention range differences lead to a divergence in the similarity of tokens that were initially similar, such as corresponding tokens from adjacent frames.
-This divergence accumulates as layers deepen.
-Based on this observation, FrameFusion prioritizes token merging at shallower layers rather than deeper layers to fully utilize the polarized high similarities.
-3.4 How Consistent Is the Ranking of Token Similarity Across Layers?
-The ranking consistency of token properties guides the compression strategy designs.
-For importance-based token pruning, the consistency across layers determines whether tokens deemed unimportant at shallow layers might become important at deeper layers.
-Early works [2, 23] adopt cascaded token pruning, where pruned tokens are no longer accessible in subsequent layers.
-These methods prioritize efficiency by eliminating token computations in both attention and Feed Forward Network (FFN) modules, and the importance is only computed once for pruned tokens.
-In contrast, more recent works [20, 17, 35] adopt non-cascaded pruning, which maintains the accessibility of all tokens across all layers by only pruning the KV-Cache at the attention layer.
-Yet, these methods do not reduce computation for the FFN and incur higher costs due to the need to compute importance for all tokens at every layer.
-Previous researches [8, 25] highlight the inconsistency of token importance across layers, which justifies re-evaluating pruning decisions at each layer.
-However, does a similar conclusion hold for token similarity?
-Figure 4: Spearman Rank Correlation between adjacent layers for the Llava-Video-7B model.
-To quantitatively assess the ranking consistency of token importance and token similarity, we calculate their respective Spearman Rank Correlations (SRC) [30] between adjacent layers.
-The SRC measures the degree to which the ranking of metrics in one layer correlates with the rankings in other layers.
-As shown in Figure 4, the SRCs for token similarity approach 1 across layers, indicating consistent ordering.
-In contrast, the SRCs for token importance are consistently lower, indicating weaker consistency across layers.
-We observe the following:
-Observation 3. Token similarity demonstrates high rank consistency across different model layers.
-Combining Observations 3 and 4, we conclude an interesting phenomenon: although the token similarity distribution condenses as layers deepen, tokens with the highest similarities remain the most similar.
-Motivated by this high consistency in similarity, FrameFusion employs similarity-based merging in a cascaded manner.
-Once highly-similar tokens are merged at shallow layers, they are not separated in subsequent computations.
+The line overlay shows the mean token similarity per layer.
+To identify the most effective layers for token compression, we analyze the distribution of token similarity values across model layers.
+Figure 3 shows that, although the mean token similarity remains relatively stable across layers, the distribution shifts noticeably:
+Observation 2. High token similarity values decrease in deeper model layers.
+This trend is particularly evident in Llava-Video model.
+While the mean similarity decreases only marginally (from 0.62 to 0.60), the distribution significantly condenses as layers deepen.
+Specifically, the 30th percentile similarity value decreases from 0.90 at the first layer to 0.72 at the last.
+Additional numerical results are in Appendix 10.1.
+The causal attention mechanism in LLMs contributes to this shift because tokens at later positions can aggregate information from earlier tokens, but not vice versa.
+This directional aggregation causes initially similar tokens, such as those corresponding across adjacent frames, to diverge increasingly at deeper layers.
+Given this observation, FrameFusion prioritizes token merging at shallower layers to effectively leverage these initially high similarities.
+3.4 Is Token Similarity Ranking Consistent Across Layers?
+The consistency of token similarity rankings across layers determines whether tokens that are similar at shallow layers remain similar in deeper layers.
+Figure 4: Spearman Rank Correlation (SRC) between adjacent layers for the Llava-Video-7B model.
+We first quantify the consistency of token similarity rankings between adjacent layers using Spearman Rank Correlation (SRC) [35], which measures the correlation of rankings across different layers.
+For comparison, we also compute the SRC for token importance.
+As shown in Figure4, token similarity maintains consistently high SRC values approaching 1, indicating stable rankings across layers.
+In contrast, token importance shows lower and unstable SRC values, suggesting that tokens considered unimportant at shallow layers might become important at deeper layers.
 Figure 5: The top-30% retention rate across model layers using different retention metrics and starting layers.
-We further validate this cascaded design choice by quantifying the top-k retention rate across model layers.
+We further examine ranking consistency between shallow and deep layers using the top-k retention rate.
 For notion simplicity, we define token uniqueness as 1 minus similarity.
-To calculate the top-k retention rate, we identify the set of top-$k$ tokens under token uniqueness or importance metrics at each layer.
-The top-k retention rate for layer $l$, starting from layer $i$, is defined as the intersection rate between the top-k tokens from both layers $l$ and $i$.
-To formalize the calculation of the top-k retention rate, we denote the set of indices of the top-$k$ retaining tokens based on the scoring function $f:\mathbb{R}^{N\times d}\rightarrow\mathbb{R}^{N}$ at layer $l$ as follows:
+To calculate the retention rate, we identify the top-$k$ tokens at each layer based on token uniqueness or importance.
+The top-k retention rate at layer $l$ with respect to a starting layer $i$ is defined as the intersection ratio between the top-k tokens at layers $l$ and $i$.
+Formally:
 $T^{(l)}=\left\{i\;\big{|}\;i\in\operatorname{Top-K}(f(\mathbf{X}^{(l)}))\right\},$ (2)
-where scoring function $f$ can be token importance or token similarity, $\operatorname{Top-K}:\mathbb{R}^{N}\rightarrow\mathbb{N}^{K}$ function returns the indices of the top-k values.
-The top-k retention rate $R^{(l)}_{i}$ of layer $l$ with the starting layer $i$ is calculated as:
-$R^{(l)}_{i}=|T^{(l)}\cap T^{(i)}|/|T^{(i)}|.$ (3)
-As shown in Figure 5, we quantify the top-30% retention rate with shallow starting layers 00 to $2$ under different metrics.
+where $f$ represents token uniqueness or token similarity, and $\operatorname{Top-K}$ selects indices of the top-k values. The retention rate $R^{(l)}i$ is then calculated as:
+$R^{(l)}{i}=|T^{(l)}\cap T^{(i)}|/|T^{(i)}|.$ (3)
+As shown in Figure 5, we calculate the top-30% retention rate at shallow starting layers 00 to $2$ under different metrics.
 The token similarity exhibits a much higher retention rate than token importance.
-This finding validates the superiority of using token similarity over importance for cascaded token pruning, achieving a better accuracy trade-off.
+Based on these analyses, we conclude:
+Observation 3. Token similarity rankings are more consistent across layers than token importance rankings.
+Combining Observations 3 and 5, we highlight a key phenomenon: while the similarity between highly similar tokens decreases at deeper layers, the relative similarity rankings of these tokens remain stable.
+Motivated by this stable ranking, FrameFusion merges highly similar tokens at shallow layers, maintaining this reduction throughout subsequent layers.
 4 FrameFusion Design
-Figure 6: FrameFusion first merges tokens with similarities above a specified threshold at shallow layers, then applies top-$k$ importance pruning to comply with the given computational constraints.
-Building upon the observations from Section 3, we introduce FrameFusion, a novel token compression technique for video LVLMs, exploring a new perspective of token similarity.
-The design of FrameFusion is detailed in Section 4.1, followed by an explanation of the rationale behind each key design choice in Section 4.2.
+Figure 6: FrameFusion first (a) merges tokens with similarities above a specified threshold at shallow layers, then (b) applies top-$k$ importance pruning to comply with the given computational constraints.
+(c) Tokens are permanently reduced for subsequent layers.
+Building upon the observations from Section 3, we propose FrameFusion, a novel token compression method for video LVLMs, exploring the new perspective of token similarity.
+The detailed design is introduced in Section 4.1, followed by the rationales behind key design choices in Section 4.2.
 4.1 Two-Stage Token Compression
-The core concept of FrameFusion is illustrated in Figure 1: unlike traditional methods that primarily employ importance-based token pruning, FrameFusion also emphasizes similarity-based token merging.
-It merges similar tokens and prunes unimportant ones, retaining only those that are both important and unique.
-The two-stage token reduction process of FrameFusion is shown in Figure 6.
+The core idea of FrameFusion is illustrated in Figure 1.
+Unlike traditional methods that solely rely on importance-based pruning, FrameFusion additionally merges similar tokens before pruning, retaining only those that are both important and unique.
+The two-stage token compression approach is depicted in Figure 6.
 Merging stage.
-In the initial merging stage, FrameFusion utilizes token similarity to merge vision tokens.
-It computes the token similarity $S^{(l)}$ at the current layer using Equation 1, where only $N$ cosine similarities are computed between the corresponding vision tokens of adjacent frames.
-Tokens that exceed the similarity threshold $S_{\text{threshold}}$ are grouped with their analogous tokens from the previous frame.
-These merging groups are transitive, allowing concatenated groups to form a larger group containing more than two tokens as shown in Figure 6.
-Within each group, FrameFusion performs element-wise averaging of all tokens, assigning the result to the first token in the group.
-This forward merging strategy ensures that subsequent vision tokens can still aggregate information from all preceding tokens using causal attention.
-The merging stage is applied across successive shallow LLM layers to progressively merge similar tokens, until the number of similar tokens falls below the threshold $N_{\text{threshold}}$ at a specific layer.
-Merging occurs right before the FFN module; however, to reduce the number of tokens for the attention module at the first LLM layer, an additional merging step is performed before it.
-After the merging stage, the remaining unique tokens advance to the pruning stage.
+In the first merging stage, FrameFusion utilizes token similarity to merge visual tokens.
+Specifically, it computes token similarity $S^{(l)}$ at each shallow layer according to Equation 1, considering only $N$ cosine similarities between corresponding tokens from adjacent frames.
+Tokens whose similarity exceeds a predefined threshold ($S_{\text{threshold}}$) are grouped with their corresponding tokens from previous frames.
+These merging groups are transitive, allowing concatenated groups to form a larger group containing more than two tokens (as shown in Figure 6).
+Within each group, FrameFusion performs element-wise averaging of all tokens and assigns the averaged result to the earliest token in the group.
+This forward merging strategy ensures that subsequent visual tokens can still aggregate information from all preceding tokens using causal attention.
+This merging procedure is progressively applied at successive shallow LLM layers until the number of similar tokens falls below a predefined threshold ($N_{\text{threshold}}$).
+Merging is applied before the feed-forward network (FFN) module.
+However, to further reduce token numbers in the attention module of the first LLM layer, an additional merging step is performed before it.
+After completing the merging stage, the remaining tokens proceed to the pruning stage.
 Pruning stage.
 After the merging stage, FrameFusion further prunes unimportant tokens.
-As discussed in Section 3.1, FrameFusion uses cumulative attention scores to represent token importance.
-Based on a user-defined computational cost budget, FrameFusion calculates the maximum number of remaining tokens $k$ that fits within the budget.
-It then applies top-$k$ importance pruning to retain only the important tokens from the remaining unique ones.
-Through the merging and pruning stages, only the unique and important vision tokens are retained for subsequent processing, significantly boosting LVLM efficiency.
+As defined in Section 3.1, FrameFusion uses cumulative attention scores to represent token importance.
+Given a user-specified computational budget, FrameFusion determines the maximum allowable number of tokens ($k$).
+It then applies pruning to retain only the top-$k$ tokens based on importance scores.
+Through the merging and pruning stages, FrameFusion effectively retains only unique and important visual tokens for subsequent processing, significantly enhancing the efficiency of LVLMs.
+We also investigate the effect of different $S_{\text{threshold}}$ and $N_{\text{threshold}}$.
+Experimental results indicate that adjustments to these thresholds result in only minor variations in model performance.
+Detailed results are presented in Section 8.3.3.
 4.2 Design Choice Rationales
-In this subsection, we explain the key design choices of FrameFusion with rationales from key observations in Section 3.
+In this subsection, we explain the rationales behind the key design choices of FrameFusion, grounded in the observations from Section 3. More empirical evidences are presented in Section 5.5.
+Design Choice 1. FrameFusion computes token similarities only between corresponding visual tokens of adjacent frames.
 Unlike token importance, which reuses the existing $N\times N$ attention scores, token similarity introduces a new, orthogonal metric.
-To avoid additional $N\times N$ similarity computations for all token pairs, we leverage Observation 2 to compute only empirically similar token pairs with an $O(N)$ complexity:
-Design Choice 1. FrameFusion computes token similarities only between corresponding vision tokens of adjacent frames.
-Another key design choice is determining the layers at which to perform token reduction.
-For importance-based pruning, previous study [2] observes the reduction in vision token importance after the initial two layers for LVLMs.
-Thus, it advises applying importance-based pruning after the second layer.
-Other works also confirm the accuracy advantage of aggressive pruning only at deeper layers [8, 1].
-In contrast, merging relies on high token similarities and therefore prefers shallow layers, according to Observation 3.
-Based on the contradictory preferences of similarity-based merging and importance-based pruning, FrameFusion is designed as follows:
-Design Choice 2. FrameFusion applies token merging at the initial successive layers, then pruning at a later layer.
-The final key design choice is whether merged tokens should remain reduced in subsequent layers, i.e., whether to use cascaded token merging.
-This choice depends on whether tokens merged in shallow layers are likely to be needed in deeper layers.
-Given the high rank consistency and retention rate of token similarities discussed in Observation 4, we adopt the following design to achieve a better efficiency trade-off:
+To avoid additional $N\times N$ similarity computations for all token pairs, we exploit Observation 2 to compute only empirically similar token pairs with an $O(N)$ complexity.
+Design Choice 2. FrameFusion applies token merging at the initial successive layers, followed by pruning at deeper layers.
+Another critical design choice involves determining the appropriate layers for different token reduction methods.
+For importance-based pruning, previous studies indicate a decline in visual token importance after the initial layers [3], recommending less pruning at shallow layers [9, 1].
+In contrast, similarity-based merging depends on initially high token similarities, preferring shallow layers as per Observation 3.
+Given these contrasting preferences, FrameFusion employs merging at shallow layers and pruning at deeper layers to optimize both similarity and importance metrics.
 Design Choice 3. FrameFusion merges tokens in a cascaded manner.
-These rationales for the three key design choices in FrameFusion enhance the explainability of its better performance.
+The final design choice addresses whether merged tokens should remain combined across subsequent layers (cascaded merging) or be individually reconsidered at each layer (non-cascaded merging).
+Specifically, as shown in Figure 6(c), cascaded merging permanently reduces the token count once tokens are merged, significantly lowering computational costs in both feed-forward network (FFN) and attention modules at subsequent layers.
+In contrast, non-cascaded merging maintains the original token count at every layer.
+It selectively reduces computations within certain modules, typically pruning only the Key and Value matrices in attention layers [26, 20, 43].
+Although non-cascaded merging retains flexibility by potentially reusing tokens at deeper layers, it incurs additional computational overhead due to repeated similarity evaluations and unchanged FFN computations.
+Given these accuracy-efficiency trade-offs, the optimal merging strategy depends on whether tokens merged at shallow layers remain similar in deeper layers.
+Considering the higher consistency of token similarity rankings across layers (Observation 5), FrameFusion employs cascaded merging to eliminate unnecessary computations.
+These rationales illustrate the motivation behind each design choice, clarifying the superior performance of FrameFusion.
 5 Experiment
 5.1 Setups
-Baselines. We compare FrameFusion with state-of-the-art token pruning baselines, StreamingLLM [26] and FastV [2]. Hyperparameters follow the official implementations and are detailed in Appendix7.
+Baselines.
+We compare FrameFusion with state-of-the-art token pruning baselines, StreamingLLM [31], FastV [3], and PruMerge [25].
+Hyperparameters adhere to respective official implementations and are detailed in Appendix 7.
+Models.
+We evaluate our approach across six video LVLMs from diverse model families and sizes, including lmms-lab models: LLaVA-Video-{7B,72B}-Qwen2 (denoted Llava-Video-{7B,72B}) [42]; NVLabs models: NVILA-Lite-2B, NVILA-8B-Video, NVILA-Lite-15B-Video (denoted NVILA-{2B,8B,15B}) [23]; and OpenBMB model MiniCPM-V-2_6 (denoted MiniCPM-V-8B) [34].
+The PruMerge baseline is incompatible with MiniCPM-V due to its Q-Former architecture and is excluded for this model.
+Benchmarks.
+We use lmms-eval [40] as the primary evaluation framework and test five video benchmarks: Video Needle In A Haystack (VideoNIAH)[44] for visual content retrieval; NExT-QA [32] for video question-answering; and VideoMME [7], EgoSchema [24] and MVBench [19] for general video understanding, highlighting spacial and temporal understanding, respectively.
 Token Budget.
-We term the token budget as the average number of tokens required for the KV-Cache at the start of the decoding stage.
-The relative token budget, denoted as $C$, is the token budget of compression methods divided by the original input length $N$.
-The relative token budget approximates the relative KV-Cache utilization for each compression method.
-We set $C=30\%$ for all token compression methods if not specified.
-Specifically, for the cascaded methods FrameFusion and FastV, the token budget equals the average token length per layer across both the prefill and decode stages.
-For StreamingLLM, the token budget equals the token length in KV-Cache at the decode stage, which also equals the attention span during the prefill stage.
-Models.
-We evaluate our approach on widely used video LVLMs, including lmms-lab models: LLaVA-Video-{7B,72B}-Qwen2 (referred to as Llava-Video-{7B,72B}), LLaVA-NeXT-Video-32B-Qwen (referred to as Llava-Video-32B) [34], and the openbmb model MiniCPM-V-2_6 (referred to as MiniCPM-V-8B) [29].
-Benchmarks.
-We use lmms-eval [32] as the primary evaluation framework to test diverse video benchmarks, including VideoMME [6] for video understanding, NExT-QA [27] for video question-answering, and Video Needle In A Haystack (VideoNIAH)[36] for visual content retrieval.
-5.2 Accuracy-Computation Trade-off
+We define token budget as the average sequence length of KV-Cache at the start of the decoding stage.
+For cascaded methods (FrameFusion, FastV, and PruMerge), it also equals the average token length per layer of the prefill stage.
+For StreamingLLM, it equals the sink size plus window size.
+The relative token budget, denoted $C$, is the token budget divided by the original input length $N$.
+Unless specified otherwise, we set $C=30\%$ for all token compression methods.
+5.2 Computation-Accuracy Trade-off
 Figure 7: The accuracy-computation trade-offs of various token compression methods, tested on Llava-Video-7B with VideoNIAH benchmark. Original* represents the original model with reduced frame rates.
-As shown in Figure 7, we explore the accuracy-computation trade-off of different token compression methods.
-The relative FLOPs is normalized by the original dense model with the video sampling rate at 1 frame per second.
-For the original variant, the relative FLOPs is altered by directly multiplying the sampling rate with the relative token budget.
-FrameFusion advances the Pareto Front in video content retrieval accuracy and computing FLOPs across varied relative token budgets.
+Figure 7 explores the computation-accuracy trade-offs of FrameFusion by varing the relative token budget.
+The x-axis shows the relative computation FLOPs, normalized to the original dense model operating at a 1 frame-per-second (fps) sampling rate.
+The y-axis shows the VideoNIAH retrieval accuracy.
+Higher accuracy at lower FLOPs (towards the top-left) indicates better trade-offs.
+The Original baseline, which directly adjusts the sampling rate of the original model, shows that FrameFusion achieves faster accuracy gains per FLOP compared to directly increasing frame rates.
+Other baselines also show significant accuracy degradation at reduced token budgets.
+In contrast, FrameFusion maintains high accuracy even at 30% computing FLOPs, greatly advancing the Pareto Front.
+We also explore the computation-accuracy trade-offs on different models and benchmarks.
+The results are detailed in Appendix 8.1.1.
 5.3 Performance
-VideoMME $\uparrow$ NExt-QA $\uparrow$ VideoNIAH $\uparrow$
-Model Method Short Medium Long Multi Choice Open End Edit Insert1 Insert2 Avg.
-Llava-Video-7B Original 75.8 61.7 52.2 83.2 32.1 90.7 50.7 88.0 66.8
-StreamingLLM 63.4 54.1 46.4 79.0 30.3 26.0 15.3 28.7 42.9
-FastV 68.4 58.0 49.6 81.1 31.2 69.3 28.7 76.7 57.9
-Ours 74.0 59.8 50.0 81.8 31.7 90.0 48.7 87.3 65.4
-MiniCPM-V-8B Original 69.1 56.6 49.8 78.9 13.8 88.7 36.7 88.7 60.3
-StreamingLLM 61.1 51.8 48.4 76.0 23.2 22.0 15.3 28.7 40.8
-FastV 67.1 53.9 49.2 78.0 14.8 82.7 26.7 71.3 55.5
-Ours 69.7 54.1 48.3 78.2 16.3 89.3 41.3 89.3 60.8
-Table 1: The overall performance comparison across different models and methods on VideoMME, NextQA, and VideoNIAH benchmarks. All token compression methods employ 30% relative token budget.
-FrameFusion outperforms state-of-the-art token compression methods across various models and benchmarks, achieving performance comparable to the dense model at a 30% token budget (Table 1).
-Additional results for additional budgets are in Appendix 8.1.
-Across VideoMME, NExT-QA, and VideoNIAH benchmarks, FrameFusion limits relative performance drops to under 3%.
-In contrast, StreamingLLM and FastV show drops of up to 13% and 7% for VideoMME and NExT-QA, and 70% and 24% for VideoNIAH.
-Further retrieval performance details are in Appendix 8.2.
+Model Size Method VideoNIAH NExT-QA VideoMME EgoSchema MVBench Avg.$\uparrow$
+edit insert1 insert2 mc oe w/o sub. w/ sub.
+Llava-Video 7B Original 90.7 50.7 88.0 83.2 32.1 63.2 69.8 53.4 61.9 65.9
+StreamingLLM 26.0 15.3 28.7 79.0 30.3 54.7 65.5 46.6 55.2 44.6
+FastV 69.3 28.7 76.7 81.1 31.2 58.7 67.0 50.1 58.0 57.9
+PruMerge 83.3 36.0 83.3 79.4 30.8 60.0 68.6 50.7 56.0 60.9
+Ours 90.0 48.7 87.3 81.8 31.7 61.3 69.9 53.0 59.7 64.8
+72B Original 89.3 66.0 88.0 85.3 32.3 70.9 77.3 65.0 63.9 70.9
+StreamingLLM 33.3 20.0 35.3 81.9 30.6 62.6 72.9 60.2 58.0 50.5
+FastV 22.0 48.7 77.3 83.7 31.5 65.9 73.7 62.6 61.7 58.6
+PruMerge 85.3 58.0 86.0 82.0 31.4 66.7 74.8 62.6 58.6 67.3
+Ours 90.0 63.3 88.0 84.6 32.0 69.0 76.7 63.2 63.0 70.0
+NVILA 2B Original 90.0 22.0 87.3 71.2 6.6 50.9 53.2 42.3 50.7 52.7
+StreamingLLM 26.0 12.7 34.7 69.0 5.8 45.7 50.1 40.7 49.1 37.1
+FastV 50.7 14.7 56.7 70.7 7.2 46.7 50.6 41.1 50.1 43.2
+PruMerge 27.3 31.3 81.3 67.7 11.1 47.3 50.4 42.2 48.0 45.2
+Ours 89.3 27.3 87.3 71.8 20.1 50.4 53.1 45.2 49.5 54.9
+8B Original 98.7 40.7 100.0 81.7 33.0 63.9 68.3 52.0 67.5 67.3
+StreamingLLM 30.0 17.3 41.3 78.4 30.8 54.3 63.7 46.2 58.1 46.7
+FastV 87.3 33.3 90.7 80.4 32.5 59.5 66.8 50.5 64.5 62.8
+PruMerge 4.7 32.0 93.3 77.1 31.4 56.9 65.1 49.4 57.9 52.0
+Ours 96.0 38.0 98.7 80.7 32.5 61.1 68.2 52.5 65.0 65.9
+15B Original 95.3 42.0 100.0 78.7 30.9 65.8 72.3 58.2 60.5 67.1
+StreamingLLM 34.0 18.7 34.0 74.0 28.5 58.5 65.1 53.7 55.0 46.8
+FastV 48.7 24.7 80.7 77.0 30.6 60.6 69.1 56.7 57.3 56.2
+PruMerge 19.3 43.3 98.0 72.4 30.0 59.3 68.4 52.3 52.8 55.1
+Ours 94.0 52.7 99.3 77.7 31.2 63.5 70.8 57.8 58.4 67.3
+MiniCPM-V 8B Original 88.7 36.7 88.7 78.9 13.8 58.5 60.3 53.4 55.0 59.3
+StreamingLLM 22.0 15.3 28.7 76.0 23.2 53.8 56.7 48.2 51.3 41.7
+FastV 82.7 26.7 71.3 78.0 14.8 56.7 58.2 51.8 53.2 54.8
+Ours 89.3 41.3 89.3 78.2 16.3 57.4 59.5 52.3 53.6 59.7
+Table 1: Performance comparison across different model families, sizes, and methods on five benchmarks at a 30% relative token budget.
+Overall Performance.
+FrameFusion consistently outperforms state-of-the-art token compression methods across multiple model families, sizes, and benchmarks, matching dense model performance at a 30% token budget (Table 1).
+FrameFusion maintains a maximum relative average performance drop of just 2.4% across six models, whereas StreamingLLM, FastV, and PruMerge incur drops of 35.8%, 19.7%, and 15.4%, respectively.
+On VideoNIAH benchmark, FrameFusion shows a maximum relative drop of only 2.8%, compared to 38.7-69.5% for other methods.
+VideoMME is sensitive because it relies on spatial and temporal details, which cannot be answered by LLM common sense alone.
+Excluding VideoNIAH, FrameFusion’s maximum drop is 3.6%, significantly lower than the 8.3-14.2% seen in baselines.
+Detailed statistics are in Appendix 8.1.
+Figure 8: The VideoNIAH performances for the Llava-Video-7B across various numbers of input frames.
+Scaling Input Length.
+Figure 8 investigates model performance across varying frame counts, from 8 to 256 frames.
+FrameFusion consistently outperforms baseline methods, achieving retrieval accuracies close to the original dense model.
+As the frame count increases, the performance gap between FrameFusion and the dense model shrinks from 4.6% to 0.5%, significantly outperforming the best baseline, which has gaps ranging from 12.8% down to 5.8%.
+We also explore the scalability of FrameFusion across different token budgets, models and benchmarks in Appendix 8.1.
 5.4 Efficiency
-Figure 8: Runtime and memory breakdown of Llava-Video-7B on a single A100-80GB GPU using FrameFusion.
-A relative token budget of 1.0 represents the original dense model.
-Numbers on bars show (a)LLM and end-to-end speedups and (b) LLM’s KV-Cache and total relative memory.
-We evaluate efficiency on a single NVIDIA A100-80GB GPU, averaging results over 128 videos from the VideoMME benchmark.
-Figure 8 highlights how FrameFusion improves runtime and memory efficiency for the colored components.
+Figure 9: End-to-end runtime and memory consumption of Llava-Video-7B and 72B models with FrameFusion, using one and four NVIDIA A100-80GB GPUs for 7B and 72B models, respectively.
+We evaluate the wall-clock speedup and GPU memory reduction of FrameFusion in Figure 9, across 7 to 72B model sizes and 32 to 256 frames.
+Results are averaged over 128 videos from the VideoMME benchmark.
+Additional results, including runtime and memory breakdowns, evaluations on more models and token budgets, and token reduction details, are provided in Appendix 8.2.
+Asymptotic complexity analysis are in Appendix 9.
 Time.
-FrameFusion achieves 3.4-4.4$\times$ speedups with a 30% token budget across various input frames.
-Including other runtimes like ViT and CPU calls, FrameFusion maintains 1.6 – 1.9$\times$ end-to-end speedups.
-Speedup results for additional model sizes are provided in Appendix 8.3.
+FrameFusion achieves end-to-end speedups of 1.6-3.6$\times$ with a 30% token budget.
+Compression stages in FrameFusion account for only 0.6-4% of the total runtime.
+The speedup grows as the number of frames increases, owing to the reduced $O((CN)^{2})$ complexity in attention computations.
+Since FrameFusion speedups the LLM computation in LVLMs, it yields greater speedups with the larger 72B model, where overhead such as video sampling and ViT encoding becomes proportionally less significant.
 Memory.
-FrameFusion reduces LLM memory consumption for KV-Cache and corresponding activations (noted as “KV-Cache” in Figure 8) to 44% – 51% with a 30% token budget, enabling more frames to fit within GPU memory limits.
-5.5 Scalability
-Scaling Model Size.
-Figure 9: The VideoMME performance for each category across Llava-Video-7B, 32B, and 72B for different methods. All scores are normalized by the original model.
-We test the scalability of FrameFusion across diverse model parameter sizes with the Llava-Video model family.
-As shown in Figure 13, FrameFusion consistently outperforms FastV baseline across all model sizes and VideoMME categories, demonstrating comparative performance with the original model at a 30% relative token budget.
-The numeric results are shown in Appendix 8.5.
-Figure 10: The VideoNIAH performance for the Llava-Video-7B across various numbers of input frames.
-Scaling Input Length.
-Figure 10 shows that FrameFusion surpass other baselines to achieve high retrieval accuracies across diverse input number of frames, demonstrating comparable performance with the original dense model.
-Figure 8 shows increasing LLM speedups and memory reductions with more input frames.
+FrameFusion reduces GPU memory for KV-Cache and activations by the factor of token budget.
+As the number of frames grows and the dominance of model parameters decreases, memory savings continue to improve.
 5.5 Ablation Study
-We validate the impact of each observation and design choice in FrameFusion, demonstrating their individual effectiveness in this section.
-Figure 11: Top-k hit numbers for different similarity calculation strategies on the initial layers of the Llava-Video-7B model.
+We validate the impact of each observation and design choice in FrameFusion, showing their individual effectiveness with average scores on VideoNIAH, VideoMME (without subtitles), and NExT-QA benchmarks.
+We further analyze the effects of similarity computation strategies, distance metrics, similarity thresholds, and positional embeddings, as detailed in Appendix 8.3.
 Design Choice 4.2.
-Instead of computing $N\times N$ token similarities with significant overhead, FrameFusion calculates $N$ token similarities only between corresponding vision tokens in adjacent frames.
-We compare our approach with two common alternatives: 1.
-The adjacent strategy, which computes $N$ similarities between adjacent image patches (i.e., adjacent vision tokens).
+Instead of computing $N\times N$ cosine similarities with significant overhead, FrameFusion calculates only $N$ token similarities between corresponding visual tokens in adjacent frames.
+We compare our approach with two common alternative strategies: 1.
+Adj.
+token, which computes $N$ similarities between adjacent image patches (i.e., adjacent visual tokens).
 2.
-The random strategy, which calculates similarities for $N$ randomly selected token pairs.
-All three strategies are compared against the posterior optimal upper bound, which merges the most similar tokens using the full $N\times N$ similarity computation.
-As shown in Figure 11, we evaluate the top-K hit number for each strategy.
-This metric indicates how many of the top-k similar token pairs identified by a given strategy align with those selected by the posterior optimal approach.
-For visualization clarity, we limit the number of frames to three (420 non-first-frame vision tokens) from each video.
-Our method consistently outperforms the adjacent and random strategies across all model layers.
-Merging Layer Budget Merging Rate Retrieval Acc.
-Original 100% 0% 76.4%
-0 50% 50.0% 76.2%
-1 50% 52.0% 76.8%
-2 50% 53.8% 76.4%
-12 50% 87.5% 74.4%
-13 50% 93.3% 64.2%
-14 50% 100.0% 48.9%
-Table 2: Performance on VideoNIAH of different merging layers with the same relative token budget of 50%.
-Method VideoMME $\uparrow$ VideoNIAH $\uparrow$
-Original 63.2 76.4
-Prune then merge 59.9 73.1
-Merge then prune 60.9 73.3
-Table 3: Performance of the Llava-Video-7B model of different orders between merging and pruning.
+Random, which calculates similarities for $N$ randomly selected token pairs.
+We compare the three similarity calculation strategies on three benchmarks at a 30% token budget.
+As shown in Table 2, our strategy achieves 11% and 8% higher average accuracies than the other two strategies.
+We also examine the choice of similarity calculation metrics by comparing cosine similarity with alternative distance metrics, including the inner product, Minkowski-2, and Minkowski-1 distance.
+Experiments show that cosine similarity yields 2-5% average score gains over the alternative metrics.
+The detailed numbers are presented in Appendix 8.3.2.
+Design 4.2 VideoNIAH VideoMME NExT-QA Avg.
+Random 60.0 59.0 56.6 58.5
+Adj. token 64.0 59.9 56.6 60.2
+Ours 76.5 61.3 56.8 64.9
+Table 2: Performance of different similarity calculation strategies with the same relative token budget of 30% on VideoNIAH, VideoMME, and NExT-QA.
+Layer Rate VideoNIAH VideoMME NExT-QA Avg
+Original 76.4 63.2 57.7 65.8
+0 50.0% 76.2 62.7 57.4 65.4
+1 52.0% 76.8 62.6 57.4 65.6
+2 53.8% 76.4 62.0 57.3 65.2
+12 87.5% 74.4 60.3 56.4 63.7
+13 93.3% 64.2 57.9 55.6 59.2
+14 100.0% 48.9 52.6 50.9 50.8
+Table 3: Performance of Llava-Video-7B with cascaded token merging at different layers, where merging rates are adjusted to maintain an average token count (token budget) of 50% across all layers.
+Design 4.2 VideoNIAH VideoMME NExT-QA Avg
+Prune $\rightarrow$ merge 73.1 59.9 55.9 63.0
+Merge $\rightarrow$ prune 73.3 60.9 56.6 63.6
+Table 4: Performance of the Llava-Video-7B model with different orders of merging and pruning.
 Design Choice 4.2.
 FrameFusion first merges at the initial layers, then prunes at subsequent layers.
-We validate the influence of the merging layer position.
-As shown in Table 2, given the same token reduction budget, merging at shallow layers shows significantly better performance than merging at deeper layers.
+We evaluate the influence of the merging layer position.
+As shown in Table 3, delaying the merging to later layers requires a higher merging rate to meet the given relative token budget of 50%.
+The reduced high similarity values at deeper layers also causes significant performance drops compared with merging at shallower layers.
 Additionally, we examine the effect of the order of merging and pruning.
-As shown in Table 3, with a fixed token budget of $C=0.3$ and the same number of tokens reduced in layers 1 and 2, merging before pruning achieves better performance compared to pruning before merging.
-[…]
-#Frame w. Cascaded Original Ours (Cascaded)
-8 92.8% 20.0 20.7
-16 90.0% 29.1 28.0
-32 81.5% 43.8 42.9
-64 72.2% 76.4 75.6
-Table 4: Equivalent relative token budgets and VideoNIAH scores of cascaded merging across different numbers of frames, aligned to the minimum FLOPs achievable by non-cascaded merging.
+As shown in Table 4, with a fixed token budget of 30% and the same number of tokens reduced in layers 1 and 2, merging before pruning achieves better performance compared to pruning before merging.
+Design 4.2 VideoNIAH VideoMME NExT-QA Avg
+Non-cascaded 74.9 60.0 55.4 63.4
+Cascaded 76.0 62.8 57.5 65.4
+Table 5: Performance of the Llava-Video-7B model with different orders of merging and pruning.
 Design Choice 4.2.
 FrameFusion adopts a cascaded merging strategy, where merged tokens remain reduced across layers to maximize efficiency.
-To evaluate the accuracy-efficiency trade-offs of cascaded merging, we use non-cascaded merging (KV-cache merging) as a baseline.
-This baseline reduces only the computation of attention matrices while leaving FFN computations unchanged.
-We calculate the minimum FLOPs achievable with non-cascaded merging by removing all KV-caches.
-As shown in Table 4, under the same FLOPs, FrameFusion requires token budgets of only 72% to 93% when using cascaded pruning—substantially more relaxed than the 30% budgets used in the main experiments.
-Across different numbers of frames, FrameFusion achieves performance comparable to the original model, while the non-cascaded model will fail due to the absence of KV-caches.
+To evaluate the accuracy-efficiency trade-offs of cascaded merging, we also implement non-cascaded merging for comparison.
+Following previous works [26, 20, 43], this baseline performs the exact same merging strategy, but only on Key Value metrics in the attention module to avoid prominently removing any tokens.
+As a result, it leaves FFN computations unchanged.
+Its merging rate across all layers are set to be 30%.
+Our cascaded counterpart uses the same computation FLOPs, translating to an 84% relative token budget.
+As shown in Table 5, under the same FLOPs, FrameFusion outperforms the non-cascaded counterparts by 3% in average score.
+showing comparable performance to the original model.
 6 Conclusion
@@
 This approach optimizes computational efficiency and memory usage, enabling accurate video understanding with significantly fewer tokens.
-Experiments across multiple benchmarks demonstrate a 70% reduction in vision tokens with minimal performance loss, achieving over 3.1$\times$ LLM speedups and 1.6$\times$ end-to-end speedups.
+Experiments across multiple benchmarks demonstrate a 70% reduction in visual tokens with minimal performance loss, achieving 1.6-3.6$\times$ end-to-end speedups.
 FrameFusion providing new insights in token similarity for LVLMs, offering an efficient and scalable solution for real-world video language applications.
+Acknowledgement
+This work was supported by National Natural Science Foundation of China (No.
+62325405, 62104128, U19B2019, U21B2031, 61832007, 62204164, 92364201), Tsinghua EE Xilinx AI Research Fund, and Beijing National Research Center for Information Science and Technology (BNRist).
+We thank all the support from Infinigence-AI.
 Supplementary Material
 7 Detailed Experiment Setup
+7.1 Method Setup
+7.1.1 Baseline Setup
 For the baselines StreamingLLM [31] and FastV [3], we follow the official implementations and set the attention sink size of StreamingLLM to 8 and $K$ in FastV to 2.
-For FrameFusion, the merging ratios across layers are controlled by two hyperparameters: $S_{\text{threshold}}$ and $N_{\text{threshold}}$, as discussed in Section 4.1.
+7.1.2 FrameFusion Setup
+Figure 10: The workflow of FrameFusion when applied to LVLMs.
+At each layer, FrameFusion performs merging, pruning, or no action between the self-attention and feed-forward layers, depending on the current stage.
+The stage initially starts as “merge” and updates according to transition conditions.
+Workflow details.
+For FrameFusion, token merging is only applied to visual tokens because they dominate input length and show higher similarity between adjacent frames, enabling $O(N)$ complexity merging.
+The detailed workflow of FrameFusion is shown in Figure 10.
+Hyperparameters. The merging ratios across layers are controlled by two hyperparameters: $S_{\text{threshold}}$ and $N_{\text{threshold}}$, as discussed in Section 4.1.
 $S_{\text{threshold}}$ defines the minimum cosine similarity required for two tokens to be considered similar and merged.
 Since similarity distributions vary across models, we set $S_{\text{threshold}}$ to match the median of similarity at the first model layer under typical input cases, such as 128 samples from the VideoMME dataset.
-For the Llava-Video series, we set $S_{\text{threshold}}=0.6$; while for MiniCPM-V, we set $S_{\text{threshold}}=0.7$.
+For the Llava-Video series, we set $S_{\text{threshold}}=0.6$; for MiniCPM-V, we set $S_{\text{threshold}}=0.7$; for NVILA-2B,8B,15B, We set $S_{\text{threshold}}=0.6,0.75,0.8$, respectively.
 $N_{\text{threshold}}$ determines the transition from merging to pruning.
@@
 Any remaining merging or pruning steps are skipped, and the model proceeds with a standard forward pass.
+7.2 Model Setup
+We follow the default frame count settings for all models, except for NVILA-Lite-2B.
+Since NVILA-Lite-2B is not specifically trained for video tasks, we set its frame count to 64.
+For the Llava-Video series and Minicpm-V, the frame count is set to 64, while for NVILA-Video-8B and NVILA-Video-15B, it is set to 256.
 8 Additional Experiment Results
-8.1 Performance Across Token Budgets
+8.1 Performance
+8.1.1 Computation-Accuracy Trade-off
+Figure 11: The accuracy-computation trade-offs of various token compression methods, tested on Llava-Video-7B with VideoMME benchmark. Original* represents the original model with reduced frame rates.
+Figure 12: The accuracy-computation trade-offs of various token compression methods, tested on NVILA-8B with VideoNIAH benchmark. Original* represents the original model with reduced frame rates.
+We further investigate the trade-off between computational cost and accuracy.
+We evaluate the Llava-Video-7B and NVILA-8B models on the VideoMME and VideoNIAH benchmark, respectively.
+The results are shown in Figure  11 and Figure  12.
+As the number of FLOPs decreases, other baseline methods exhibit a noticeable decline in accuracy, whereas FrameFusion maintains superior performance.
+8.1.2 Performance Across Different Input Length
+Figure 13: The VideoMME performances for the Llava-Video-7B across various numbers of input frames.
+Figure 13 presents the performance of Llava-Video-7B on the VideoMME benchmark as the number of input frames varies from 8 to 128.
+Across all configurations, FrameFusion consistently outperforms the baseline methods, demonstrating its robustness to different input length.
+8.1.3 Performance Across Different Token Budgets
 VideoMME NExt-QA-MC NExt-QA-OE
@@
 As the budget increases to 0.5 and 0.7, the maximum drops further decrease to $\leq$1.2%.
-8.2 Retrieval Analysis
-Figure 12: VideoNIAH retrieval accuracy of the Llava-Video-7B and MiniCPM-V-8B models using different token compression methods across varying video lengths and retrieval positions.
-All token compression methods employ 30% relative token budget.
-We further investigate the retrieval accuracy details with the VideoNIAH benchmark, as shown in Figure 12.
-FrameFusion demonstrates similar retrieval performance as the original dense model, with consistent performance across lengths and positions.
-In contrast, StreamingLLM hardly retrieves the initial frames of the video.
-FastV does not show particular failure patterns but undergoes uniform performance degradation across grids.
-8.3 Speedup Across Model Sizes
-We evaluate the scalability of FrameFusion ’s speedup across different model sizes, as shown in Figure 13.
-To accommodate the increased KV-Cache and memory overhead, we distribute the models across multiple GPUs.
-With larger models, FrameFusion achieves greater end-to-end speedups, delivering 2.8×2.8$\times$2.8 × for Llava-Video-32B on two GPUs and 3.2×3.2$\times$3.2 × for Llava-Video-72B on four GPUs at a 30% token budget.
-Figure 13: Runtime of (a) Llava-Video-32B on two A100-80GB GPUs and (b) Llava-Video-72B on four A100-80GB GPUs using FrameFusion. Numbers on bars indicate LLM and end-to-end speedups.
-8.4 Token Reduction Details
-Figure 14: Average number of tokens per layer in the Llava-Video-7B model with FrameFusion at different relative token budgets. Error bars represent variance across data items.
-FrameFusion reduces computational cost through both token merging and pruning.
-Using 128 samples from the VideoMME dataset with the Llava-Video-7B model, we calculate the token count per layer.
-As shown in Figure 14, FrameFusion progressively reduces tokens per layer, achieving the desired relative token budget (represented by the area under the line).
-8.5 Scalability Numeric Results
+8.1.4 Performance Across Different Models
 We present the detailed numeric results of the scalability experiments in Section 5.3.
-Model Method Knowledge Film & Television Sports Competition Artistic Performance Life Record Multilingual
-Llava-Video-7B Original 63.1 67.2 61.8 61.7 63.7 58.9
-StreamingLLM 55.1 57.2 56.0 54.2 52.9 48.9
-FastV 59.1 60.0 58.9 57.8 58.1 55.6
-Ours 62.7 63.6 58.0 61.7 60.8 56.7
-Llava-Video-32B Original 60.9 68.9 57.3 59.2 58.1 58.9
-StreamingLLM 53.1 55.6 50.2 53.3 49.5 53.3
-FastV 56.1 62.2 55.6 56.7 54.3 50.0
-Ours 58.9 65.6 55.6 59.2 54.3 54.4
-Llava-Video-72B Original 73.2 74.4 68.0 71.4 68.9 62.2
-StreamingLLM 65.7 66.7 59.3 65.6 58.7 58.9
-FastV 66.8 72.8 61.1 69.2 63.2 61.1
-Ours 72.2 73.1 65.6 69.7 65.9 61.1
+Figure 14: The VideoMME performance for each category across Llava-Video-7B, 32B, and 72B for different methods. All scores are normalized by the original model.
+Model Method Short Medium Long KL FT SC AP LR ML
+Llava-Video-7B Original 75.8 61.7 52.2 63.1 67.2 61.8 61.7 63.7 58.9
+StreamingLLM 63.4 54.1 46.4 55.1 57.2 56.0 54.2 52.9 48.9
+FastV 68.4 58.0 49.6 59.1 60.0 58.9 57.8 58.1 55.6
+PruMerge 69.7 60.1 50.2 59.1 63.6 59.1 58.9 60.6 57.8
+Ours 74.0 59.8 50.0 62.7 63.6 58.0 61.7 60.8 56.7
+Llava-Video-72B Original 80.9 69.7 62.1 73.2 74.4 68.0 71.4 68.9 62.2
+StreamingLLM 68.2 59.9 59.8 65.7 66.7 59.3 65.6 58.7 58.9
+FastV 73.0 64.9 60.2 66.8 72.8 61.1 69.2 63.2 61.1
+PruMerge 74.0 65.8 60.3 70.4 73.6 62.9 68.3 61.6 54.4
+Ours 78.3 67.9 60.9 72.2 73.1 65.6 69.7 65.9 61.1
+NVILA-2B Original 61.4 48.9 42.4 47.2 56.4 49.8 55.0 51.0 52.2
+StreamingLLM 52.9 43.9 40.3 43.0 49.2 46.2 50.6 44.4 43.3
+FastV 53.7 45.6 40.8 43.8 49.7 46.2 51.4 46.7 43.3
+PruMerge 53.9 45.0 43.1 43.5 52.8 45.8 51.7 48.1 45.6
+Ours 61.3 47.0 43.0 48.3 55.6 48.2 55.3 49.8 45.6
+NVILA-8B Original 74.9 62.1 54.7 64.8 66.4 62.2 61.9 63.7 63.3
+StreamingLLM 61.2 53.8 48.0 54.9 57.8 54.7 52.5 52.2 55.6
+FastV 72.0 56.7 50.0 60.7 62.8 57.6 57.8 58.9 57.8
+PruMerge 67.6 54.9 48.3 57.3 61.1 56.0 54.7 56.2 55.6
+Ours 74.2 57.7 51.3 60.7 65.3 59.3 58.6 62.1 58.9
+NVILA-15B Original 77.3 64.7 55.3 67.2 68.1 62.7 63.3 66.2 66.7
+StreamingLLM 63.8 57.4 54.3 60.6 60.6 55.6 58.6 56.5 60.0
+FastV 69.2 58.7 53.9 62.8 63.3 57.1 60.8 58.1 63.3
+PruMerge 66.0 59.3 52.6 61.0 61.1 55.1 57.5 59.8 61.1
+Ours 73.2 62.3 55.0 64.6 68.1 60.9 61.1 62.5 65.6
+MiniCPM-V-8B Original 69.1 56.6 49.8 59.0 63.6 54.2 63.3 54.9 60.0
+StreamingLLM 61.1 51.8 48.4 54.6 58.1 52.2 56.4 49.7 55.6
+FastV 67.1 53.9 49.2 57.2 59.2 53.8 60.8 54.6 56.7
+Ours 69.7 54.1 48.3 57.9 63.1 53.8 60.3 54.4 56.7
 Table 7: Numeric VideoMME scores of different methods and model sizes across various video categories.
-Table 6 provides the VideoMME scores for various model sizes across different video categories, offering a numerical breakdown of Figure 13.
-FrameFusion consistently outperforms the baseline methods across diverse model sizes and video categories.
+“KL”, “FT”, “SC”, “AP”, “LR”, “ML” are short for “Knowledge”, “Film & Television”, “Sports Competition”, “Artistic Performance”, “Life Record”, and “Multilingual”.
+As shown in Figure 14, FrameFusion consistently outperforms FastV baseline across all model sizes and VideoMME categories, demonstrating comparative performance with the original model at a 30% relative token budget.
+Note that the model Llava-Video-32B has been removed by its author team.
+However, in order to demonstrate the generalization capability of our FrameFusion method across variable model sizes, we still include this model in the performance and efficiency tests here.
+Table 7 provides the VideoMME scores for various model sizes across different video lengths and categories, offering a numerical breakdown of Figure 14.
 Number of frames Max.
@@
 In contrast, both StreamingLLM and FastV exhibit noticeable drops in accuracy.
+Setting Qwen2-VL-7B InternVL-2.5-8B
+Original Ours Original Ours
+w/o sub 55.9 58.4 63.1 62.3
+w/sub 60.6 61.1 66.3 64.2
+Table 9: Performance comparison between original and Framefusion of Qwen2-VL-7B and InternVL-2.5-8B on VideoMME.
+We further test the performance of two extra models: Qwen2-VL-7B and InternVL-2.5-8B. As shown in Table 9, our method performs well compared to original models on VideoMME.
+8.1.5 Retrieval Benchmark Details
+Figure 15: VideoNIAH retrieval accuracy of the Llava-Video-7B and MiniCPM-V-8B models using different token compression methods across varying video lengths and retrieval positions.
+All token compression methods employ 30% relative token budget.
+We further investigate the retrieval accuracy details with the VideoNIAH benchmark, as shown in Figure 15.
+FrameFusion demonstrates similar retrieval performance as the original dense model, with consistent performance across lengths and positions.
+In contrast, StreamingLLM hardly retrieves the initial frames of the video.
+FastV does not show particular failure patterns but undergoes uniform performance degradation across grids.
+8.1.6 Performance on Image Benchmark
+We further investigate our method’s performance on an image benchmark: MMMU-Pro-standard.
+As shown in Table 10, although our method is not designed for image inputs, it still demonstrates comparative performance.
+Model Original Fastv Ours
+NVILA-2B 23.6 23.8 23.1
+NVILA-8B 30.3 28.7 29.0
+NVILA-15B 36.1 30.6 32.8
+Table 10: The MMMU-Pro-standard performance across NVILA-2B, 8B, and 15B for different methods.
+8.2 Efficiency
+8.2.1 Efficiency Across Different Model Sizes
+We evaluate the scalability of FrameFusion ’s efficiency across different model sizes, as shown in Figure 17 and 18.
+To accommodate the increased KV-Cache and memory overhead, we distribute models across multiple GPUs.
+With larger models, FrameFusion achieves greater end-to-end speedups, delivering 2.8×2.8$\times$2.8 × for Llava-Video-32B on two GPUs and 3.2×3.2$\times$3.2 × for Llava-Video-72B on four GPUs at a 30% token budget.
+Besides, FrameFusion reduces memory consumption for KV-Cache to 37% for Llava-Video-32B and 51% for Llava-Video-72B with a 30% token budget.
+Figure 16: Runtime and memory breakdown of Llava-Video-7B on a single A100-80GB GPU using FrameFusion.
+A relative token budget of 1.0 represents the original dense model.
+Numbers on bars show (a) LLM and end-to-end speedups and (b) LLM’s KV-Cache and total relative memory.
+Figure 17: Runtime and memory breakdown of Llava-Video-32B on two A100-80GB GPUs using FrameFusion.
+A relative token budget of 1.0 represents the original dense model.
+Numbers on bars show (a) LLM and end-to-end speedups and (b) LLM’s KV-Cache and total relative memory.
+Figure 18: Runtime and memory breakdown of Llava-Video-72B on four A100-80GB GPUs using FrameFusion.
+A relative token budget of 1.0 represents the original dense model.
+Numbers on bars show (a) LLM and end-to-end speedups and (b) LLM’s KV-Cache and total relative memory.
+8.2.2 Token Reduction Details
+Figure 19: Average number of tokens per layer in the Llava-Video-7B model with FrameFusion at different relative token budgets. Error bars represent variance across data items.
+FrameFusion reduces computational cost through both token merging and pruning.
+Using 128 samples from the VideoMME dataset with the Llava-Video-7B model, we calculate the token count per layer.
+As shown in Figure 19, FrameFusion progressively reduces tokens per layer, achieving the desired relative token budget (represented by the area under the line).
+8.3 Ablation Study
+8.3.1 Similarity Computation Strategy
+Figure 20: The average token similarity of the merged tokens for the first layer of Llava-Video-7B model across various merging rates.
+We empirically study whether our approach successfully finds the most similar token pairs.
+All three $O(N)$ complexity strategies are compared against the posterior optimal upper bound, which merges the most similar tokens using the full $N\times N$ similarity computation.
+As shown in Figure 20, given different merging rate, the token pairs found by our method constantly shows the highest average similarity.
+We successfully reach 90% average similarity with only $1/10^{4}$ computing overhead.
+Further ablations are detailed in Section 5.5
+8.3.2 Distance Metrics
+FrameFusion adopts cosine similarity as the distance metric between tokens.
+To evaluate the impact of different distance metrics, we replace cosine similarity with the inner product, Minkowski-2, and Minkowski-1 distance.
+We test the performance of FrameFusion at a 30% token budget.
+As shown in Table 11, the average accuracy using cosine similarity is 2.8%, 1.4%, and 1.4% higher than the baseline metrics, respectively.
+Choice VideoNIAH VideoMME NExt-QA Avg.
+inner product 71.3 58.9 55.0 61.7
+minkowski-2 71.3 60.9 57.1 63.1
+minkowski-1 71.3 61.0 57.0 63.1
+cosine similarity 75.1 61.4 56.9 64.5
+Table 11: Performance of different distance calculation strategies with the same relative token budget of 30% on VideoNIAH, VideoMME, and NExt-QA.
+8.3.3 Choice of Similarity Threshold
+We conduct ablation studies on the sensitivity of the similarity ($S_{\text{threshold}}$) and merging-pruning transition ($N_{\text{threshold}}$) thresholds on NVILA-8B.
+As shown in Table 12, our method shows robust performance to threshold variations.
+Threshold Value VideoNIAH VideoMME NeXT-QA-mc
+$S_{\text{threshold}}$ 0.6 73.6 56.9 56.5
+0.7 (default) 73.3 57.4 56.3
+0.8 72.9 57.6 56.5
+0.9 72.2 57.7 56.3
+$N_{\text{threshold}}$ 0.1 (default) 73.3 57.4 56.3
+0.2 74.0 57.6 56.5
+0.3 74.2 57.5 56.5
+Table 12: Performance of different similarity and merging-pruning transition thresholds on VideoNIAH, VideoMME, and NExt-QA.
+8.3.4 Effect of Positional Embedding
+We investigate the impact of positional embeddings on token similarity.
+Specifically, we compare models with and without positional embedding at the first layer and analyze the resulting changes in the similarity of the input hidden states to the second layer.
+The results show that the L1-norm of the similarity matrix changes by an absolute amount of $0.0087\pm 0.0010$, corresponding to a relative change of $2.73\%\pm 0.66\%$.
+It shows that the token contents, rather than the positional embeddings, dominate token similarity.
 9 Asymptotic Complexity Analysis
@@
 Given a model with $L$ layers and a specified relative token budget $C$, FrameFusion operates in the merging stage from layer 0 to layer $K-1$, then transitions to the pruning stage at layer $K$.
-Let $N_{l}$ denote the number of tokens layer $l$ before token reduction at this layer.
+Let $N_{l}$ denote the number of tokens in layer $l$ before token reduction at this layer.
 Note that $N_{l+1}$ represents the number of tokens of layer $l$ after token reduction, and we let $N_{-1}$ equal the original input token length $N$.
@@
 $F^{\prime}(N_{l})=3N_{l}D$ (6)
-Note that the additional computation $F^{\prime}$ introduced by FrameFusion shows neglectable asymptotic complexity with respect to input length and model size, comparing with the $O(N^{2}D)$ and $O(ND^{2})$ complexities of the original model.
+Note that the additional computation $F^{\prime}$ introduced by FrameFusion shows negligible asymptotic complexity with respect to input length and model size, compared with the $O(N^{2}D)$ and $O(ND^{2})$ complexities of the original model.
 10 Additional Observation Details
@@
 As shown in Figure 21, the similarity variance decreases in the deeper layers of the model, validating Observation 3.
+No significant outliers are observed in token similarity, in contrast to the common outliers seen with respect to the magnitude of hidden features [38, 36].
 10.2 Observations on Additional Models
-In addition to the analysis of the Llava-Video model in Section 3, we conduct a similar study on the MiniCPM architecture. Results are presented in Figures 16, 17, 18, and 19.
-Overall, the conclusions align with those of the Llava-Video model, with a few notable differences: Firstly, as shown in Figure 16, MiniCPM, which incorporates Q-Former [29, 15], exhibits additional high similarity among vision tokens within the same frame.
+In addition to the analysis of the Llava-Video model in Section 3, we conduct a similar study on the MiniCPM architecture. Results are presented in Figures 22, 23, 24, and 25.
+Overall, the conclusions align with those of the Llava-Video model, with a few notable differences: Firstly, as shown in Figure 22, MiniCPM, which incorporates Q-Former [34, 17], exhibits additional high similarity among visual tokens within the same frame.
 However, the prominent 210th sub-diagonal persists, supporting our token similarity calculation strategy.
@@
 The line overlay shows the average token similarity across layers.
-Figure 18: Spearman Rank Correlation between adjacent layers for the MiniCPM-V-8B model.
+Figure 24: Spearman Rank Correlation (SRC) between adjacent layers for the MiniCPM-V-8B model.
 Figure 25: The Top-30% retention rate across model layers for the MiniCPM-V-8B model, using different retention metrics and reference layers.
@@
 Furthermore, we replace the blank regions with the average of the merged patches, and the resulting visualization is shown in Figure 28.
-As shown in the examples, FrameFusion token merging strategy successfully merges similar vision tokens, reducing the computational costs, while maintaining high vadility of the video.
+As shown in the examples, FrameFusion token merging strategy successfully merges similar visual tokens, reducing the computational costs, while maintaining high validity of the video.
 10.4 Importance-Similarity Joint-Distribution
@@
 Figure 29: Importance-similarity joint-distribution of different layers, with color intensity denoting distribution frequency.
+11 Additional Discussion on Related Works
+Prior works have also explored token merging in image-based tasks [13, 45, 25].
+For instance, while FrameFusion adopts an $O(N)$ temporal merging strategy, EVLGen [13] performs $O(N^{2})$ spatial merging via bipartite matching among tokens.
+AIM [45] similarly adopts bipartite-matching-based merging prior to the first layer of the LLM, followed by a token pruning process in subsequent LLM layers, ultimately reducing the number of visual tokens to zero.
+LLaVA-Prumerge [25] first prunes tokens at the output of the visual encoder and then merges the pruned tokens into the top-$k$ most similar remaining tokens.
+In all these methods, the similarity computation incurs a complexity of $O(N^{2})$.
+Although the computational efficiency of is comparable at the image scale ($N\approx 256$), our method scales more effectively to video scenarios where $N$ can reach 10K to 1M tokens.
 12 Limitation and Future Works
-While FrameFusion demonstrates significant improvements in token reduction and efficiency for video LVLMs, certain challenges remain for future works.
+While FrameFusion demonstrates significant improvements in token reduction and efficiency for video LVLMs, certain challenges remain for future work.
 First, the similarity-based merging process can be further refined to better handle highly diverse or complex video content, minimizing potential information loss.